### v7

# New possibilities in identification of binary choice models with fixed effects

Abstract
We study the identification of binary choice models with fixed effects. We propose a condition called sign saturation and show that this condition is sufficient for identifying the model. In particular, this condition can guarantee identification even when all the regressors are bounded, including multiple discrete regressors. We also establish that without this condition, the model is not identified unless the error distribution belongs to a special class. Moreover, we show that sign saturation is also essential for identifying the sign of treatment effects. Finally, we introduce a measure for sign saturation and develop tools for its estimation and inference.
Key words: identification, panel model, binary choice, fixed effects
1 Introduction
This paper considers panel models with binary outcomes in the presence of fixed effects. These models are convenient in economic analysis as they allow for fairly general unobserved individual heterogeneity. The nonlinear nature of the binary choice models makes it difficult to eliminate the fixed effects by differencing. Since we view the fixed effects or their conditional distribution as a nuisance parameter, the identification of the parameter of interest becomes tricky. In this paper, we provide new results and insights on what drives the identification of the coefficients on the regressors and what this means for applied work.
Consider independent and identically distributed (i.i.d) observations $\{(Y_{i},X_{i})\}_{i=1}^{n}$ with $Y_{i}=(Y_{i,0},Y_{i,1})$ and $X_{i}=(X_{i,0},X_{i,1})$ from the following model
$Y_{i,t}=\mathbf{1}\{X_{i,t}^{\prime}\beta+\alpha_{i}\geq u_{i,t}\}\qquad t\in\{0,1\},$
where the fixed effects are represented by the scalar variable $\alpha_{i}$ and $\beta$ is a non-random vector of coefficients. This is a semiparametric model as the distribution of $\alpha_{i}$ given $X_{i}$ is unrestricted. For notational simplicity, we drop the $i$ subscript in the rest of the paper. Therefore, we write $Y=(Y_{0},Y_{1})$ and $X=(X_{0},X_{1})\in\mathcal{X}_{0}\times\mathcal{X}_{1}$ with
$Y_{t}=\mathbf{1}\{X_{t}^{\prime}\beta+\alpha\geq u_{t}\}\qquad t\in\{0,1\}.$ (1)
In this paper, we mainly focus on the identification of $\beta$ but we will also discuss quantities related to treatment effects. There are roughly two approaches to identifying $\beta$, depending on whether or not we impose a parametric model on the distribution of $u_{t}$ given $(X,\alpha)$. Methods without parametric assumptions on the error distribution are typically based on Manski, (1987) and maximum-score-type estimation. This approach only assumes that the distribution $u_{t}\mid(X,\alpha)$ does not depend on $t$. In contrast, the more parametric approach relies on additional assumptions on the functional form of the distribution $u_{t}\mid(X,\alpha)$. For example, perhaps the most popular parametric assumption is that $u_{t}$’s are i.i.d logistic errors across $t$ and are independent of $(X,\alpha)$. This approach typically adopts an estimation scheme based on the (conditional) likelihood.
Despite the strong restrictions on the functional form, the approach relying on parametric restrictions has received considerable attention arguably due to identification reasons. The literature has pointed out the widespread identification failure, such as Arellano and Bonhomme, (2011). In particular, identification seems infeasible outside the logistic case unless the support of $X$ is unbounded. For example, Assumption 2 in Manski, (1987) requires the unboundedness of at least one component of $W=X_{1}-X_{0}$: for $W=(W_{1},...,W_{K})^{\prime}$ and $\beta=(\beta_{1},...,\beta_{K})^{\prime}$, there exists $k$ such that $\beta_{k}\neq 0$ and the conditional distribution $W_{k}\mid(W_{1},...,W_{k-1},W_{k+1},...,W_{K})$ has support equal to $\mathbb{R}$ almost surely. Theorem 1 of Chamberlain, (2010) goes even further: for bounded $X$, the identification fails in certain regions of the parameter space if no further restrictions are imposed on the distribution of the error term $u_{t}$.
In this paper, we provide a more precise picture than Chamberlain, (2010) by showing that identification without parametric assumptions on the error distribution is quite possible even for bounded $X$. The main motivation of this paper is to explore new possibilities without functional-form assumptions. This raises concerns about stability. If identification crucially hinges on the imposed functional form, the reliability of the analysis could be a concern. After all, if the model parameter is unidentified under every distribution function other than the logistic one, how much should we trust this model? Hence, it is helpful to explore a more robust setting. In this direction, this paper considers the identification issue and leaves the estimation problem to future research. We make the following contributions.
First, we provide tight and simple identification conditions without parametric restrictions on the error distribution. We show that a sufficient condition for identification is what we refer to as sign saturation (Assumption 2), which states that $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)$ are both strictly positive. This condition can hold with bounded regressors $X$ and allow for multiple discrete regressors and interaction terms. Moreover, we show that this is also a necessary condition for identification unless the distribution of $u_{t}$ is in a special class. Therefore, although we know (from Chamberlain, (2010)) that for bounded regressors the identification fails at some values of $\beta$, our results pinpoint these values: the identification fails exactly at points where the sign saturation fails. Therefore, whether the regressors are bounded is not what really drives the identification. The identifiability is more closely related to the sign saturation condition, which is simple and intuitive.
A key advantage of the proposed sign saturation condition is that it is stated in terms of observed variables only and does not involve the parameter that we are trying to identify. Hence, it can be used to answer the question of whether identification holds under the distribution of observed data, thereby making identification directly testable. This is different from the Manski, (1987)-type condition. For example, suppose that $K=2$, $W_{1}$ is bounded and the conditional distribution $W_{2}\mid W_{1}$ has support $\mathbb{R}$ almost surely. The identification condition in Manski, (1987) becomes $\beta_{2}\neq 0$. However, how can we check $\beta_{2}\neq 0$ when the identifiability of $\beta$ is not yet established? It does not seem obvious how to do so with the data. In contrast, under the sign saturation condition, we check statements only on observed data: $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)$.
Second, we show that the sign saturation condition is also sufficient and necessary for learning the sign of the marginal effects. Even though the magnitude of different notions of the marginal effects is usually not point identified, the sign of the effects is typically the same as the sign of a component of $\beta$. It turns out that outside a special class of distributions, if the sign saturation condition fails, we cannot guarantee to distinguish between zero effects and strictly positive effects. In many empirical studies, one important task is to check whether the identified set for the effects includes zero. Therefore, it is worthwhile to check the sign saturation condition before exploring additional assumptions that give bounds to treatment effects.
Third, we introduce a tool that can be used to assess the sign saturation condition empirically. We show that $E\max\{P(Y_{1}-Y_{0}\mid X),0\}$ and $E\min\{P(Y_{1}-Y_{0}\mid X),0\}$ can be directly estimated from the data and confidence intervals can be constructed by a simple bootstrap without any tuning parameters such as bandwidth or number of basis functions. For bounded regressors, the existing literature highlights the risk of identification failure: the identification possibly (not definitely) fails. This generic warning may have discouraged the application of fixed effect models in practice. The proposed tool serves as a more accurate diagnosis so that the identifiability can be assessed from the data.
Our work is closely related to the literature of binary response models with logistic errors. Rasch, (1960), Andersen, (1970) and Chamberlain, (1980) consider the estimation and inference in this case. The logistic link function is also singled out in Chamberlain, (2010) as the “nice” link function: the identification does not require unbounded regressors. Recently, the work of Davezies et al., (2023) points out that if there are more than 2 time periods, then the nice link function can be extended to a generalized version of the logistic distribution. This suggests that imposing a special class of parametric structures is a useful strategy of achieving identification and this special class seems to be the logistic functions. We provide new insight on this. We show that this special class also includes those such that $\dot{G}(\cdot)$ is periodic, where $\dot{G}(\cdot)$ is the derivative of $\ln\frac{F(\cdot)}{1-F(\cdot)}$ and $F(\cdot)$ is the distribution function of $u_{t}$. If $F(\cdot)$ is logistic, then $\dot{G}(\cdot)$ is a constant function, which is clearly periodic. However, we find that identification is possible for any periodic $\dot{G}(\cdot)$. Although Chamberlain, (2010) tells us that for non-logistic distributions, identification fails in an open neighborhood, we show that for any periodic $\dot{G}(\cdot)$, identification must hold in another open neighborhood. Therefore, it seems to us that the truly problematic distributions are those with non-periodic $\dot{G}(\cdot)$. Indeed, outside this special class of periodic $\dot{G}(\cdot)$, we show that identification is impossible at every point if the sign saturation condition is not satisfied. Thus, for generic distribution functions, sign saturation is a necessary condition for identification. We summarize the relation between identifiability and error distribution in Table 1.
Our work is also closely related to works that do not assume logistic errors. First, some results in the literature allow for bounded regressors but require all regressors to be continuous. For example, Assumption 3.3 of Shi et al., (2018) gives a sufficient condition for identification under bounded support. As commented in the paper, their assumption essentially requires all regressors to be continuous. Continuous distribution on all the regressors are also required by Assumption 6 of Toth, (2017) and Assumption 4’ of Gao et al., (2023). Examples on nonseparable models include Hoderlein and White, (2012), Chernozhukov et al., (2015) and Chernozhukov et al., (2019); their arguments rely on the derivatives with respect to $X$ and thus $X$ needs to be continuous. Here, we allow for one or multiple discrete regressors, which are important in empirical studies as many treatment variables are binary. Second, Corollary 4.1 of Horowitz, (2009) considers a related model in the cross-sectional setting. It is possible to translate this to the panel data in terms of conditional densities, but these are still not the most general conditions; for example, some strictly increasing continuous distribution functions have zero density almost everywhere, e.g., Salem, (1943) and Takács, (1978). Most importantly, none of the aforementioned works show whether their conditions are necessary. We contribute to the literature by finding what must be assumed and developing empirical tools to assess it.
Our necessity results complement those in Pakes and Porter, (2024). By Proposition 3 therein, $\arg\max_{\beta}E(Y_{1}-Y_{0})\cdot\mathbf{1}\{W^{\prime}\beta\geq 0\}$, the set of maximizers of the maximum score criterion function, is the sharp identification set when we only assume the stationarity of the distribution $u_{t}\mid(X,\alpha)$. Our results show that this sharp identified reduces to a singleton up to scale under the sign saturation condition. A natural question is the necessity of this condition, especially if we are willing to impose additional assumptions, namely i.i.d errors that are independent of $(X,\alpha)$. By classical results, we know that under logistic errors, $\arg\max_{\beta}E(Y_{1}-Y_{0})\cdot\mathbf{1}\{W^{\prime}\beta\geq 0\}$ is not the sharp identified set. Our results in Section 2.3 further complete the puzzle by characterizing point identification as the periodicity of $\dot{G}(\cdot)$.
2 Identification of coefficients
In the setting of (1), Manski, (1987) identifies $\beta$ with the distributional stationarity condition:
Assumption 1.
The distribution of $u_{t}\mid(X,\alpha)$ does not depend on $t\in\{0,1\}$. Let the cumulative distribution function (c.d.f) of this common distribution be $F(\cdot|x,\alpha)$. Assume that for any $(x,\alpha)$, $F(\cdot\mid x,\alpha)$ is a strictly increasing function on $\mathbb{R}$.
Assumption 1 rules out dynamic models. In this paper, we focus on static models with two time periods. Without any restriction on the magnitude of $\alpha$ and $u_{t}$ in (1), it is impossible to identify the magnitude of $\beta$ but the following result by Manski, (1987) gives the identification of the “direction” of $\beta$; in other words, $\beta$ is identified up to scaling.
Lemma 1 (Manski, (1987)).
Let Assumption 1 hold. Then with probability one,
${\rm sgn}\left(E(Y_{1}-Y_{0}\mid X)\right)={\rm sgn}\left(W^{\prime}\beta\right),$ (2)
where $W=X_{1}-X_{0}$ and ${\rm sgn}(\cdot)$ is defined as ${\rm sgn}(t)=1$ for $t>0$, ${\rm sgn}(t)=-1$ for $t<0$ and ${\rm sgn}(0)=0$.
The key idea of our result is based on the condition in (2). Clearly, the conditional mean function $E(Y_{1}-Y_{0}\mid X)$ is identified. For simplicity, suppose that $P(E(Y_{1}-Y_{0}\mid X)=0)=0$. Then (2) implies that there is a hyperplane of $W$ that perfectly classifies ${\rm sgn}(E(Y_{1}-Y_{0}\mid X))$. This is an ideal support vector machine (SVM) as illustrated in Figure 1, where the red and blue dots represent $-1$ and $1$, respectively for ${\rm sgn}(E(Y_{1}-Y_{0}\mid X))$. [^1] In Komarova, (2013), the geometry based on SVM is also to study binary response models with a median restriction; the model there does not have a panel structure. We thank Christopher Walker for this reference. If we want to identify the classification boundary in the SVM, we obviously need to have both classes (red and blue) in the data. Since these two classes correspond to $E(Y_{1}-Y_{0}\mid X)>0$ and $E(Y_{1}-Y_{0}\mid X)<0$, this simple requirement is formalized as the following sign saturation condition.
Figure 1: Lemma 1 as a support vector machine
This demonstrates the case with $W\in\mathbb{R}^{2}$. The red dots correspond to ${\rm sgn}(W^{\prime}\beta)=1$ and the blue ones correspond to ${\rm sgn}(W^{\prime}\beta)=-1$.
Assumption 2.
Both $P(E(Y_{1}-Y_{0}\mid X)>0)$ and $P(E(Y_{1}-Y_{0}\mid X)<0)$ are strictly positive.
From Figure 1, it is also clear that in addition to having both classes, we also need the data points to be “dense” around the hyperplane. For simplicity, we will require that the support of $W$ is convex and has no-empty interior so the support of $W$ is dense. This is all we need for uniquely determining the hyperplane and there is nothing about unbounded supports. We will formalize this intuition in in the next subsection and extend the results to the case of multiple discrete regressors in Section 2.2.
Remark 1.
Phrasing the problem in terms of support vector machines makes it easy to see that when all the regressors are discrete, we cannot in general expect to point identify $\beta$ up to scale. To see this, notice that by Pakes and Porter, (2024), the sharp identified set for $\beta$ is $\arg\max_{b}E(Y_{1}-Y_{0}){\rm sgn}(W^{\prime}b)$, which can be written as
$\arg\max_{b}E\left[\left|E(Y_{1}-Y_{0}\mid X)\right|\cdot{\rm sgn}(W^{\prime}\beta)\cdot{\rm sgn}(W^{\prime}b)\right].$
Therefore, if ${\rm sgn}(W^{\prime}\tilde{\beta})={\rm sgn}(W^{\prime}\beta)$ with probability one, then $\tilde{\beta}$ is in the identified set. When the regressors are discrete, we can in general move $\beta$ a bit and retain the same ${\rm sgn}(W^{\prime}\beta)$. As illustrated in Figure 2, the black and gray lines both perfectly classify the red and blue dots and thus represent two points in the identified set that are not scalar multiple of each other.
Figure 2: Discrete regressors as a support vector machine
This demonstrates the case in which $W\in\mathbb{R}^{2}$ is discrete. The red dots correspond to ${\rm sgn}(W^{\prime}\beta)=1$ and the blue ones correspond to ${\rm sgn}(W^{\prime}\beta)=-1$.
2.1 The sign saturation condition is sufficient for identification
Following Chamberlain, (2010), we assume that one of the regressors is binary in that it is equal to zero at $t=0$ and is equal to one at $t=1$; in other words, one component of $W=X_{1}-X_{0}$ is one. Thus, without loss of generality, we can partition $W=(Z^{\prime},1)^{\prime}\in\mathbb{R}^{K}$. If all the regressors are continuous, then the result can still be applied because the coefficient for the binary regressor is allowed to be zero, see Corollary 1. The first main result of this paper is to show that the sign saturation condition (together with additional weak assumptions) is enough to guarantee identification up to scaling. To state the formal result, we first recall the definition of the support of a random variable (or its corresponding probability measure): the support of probability measure $\lambda$ is the smallest closed set $A$ such that $\lambda(A)=1$, e.g., page 227 of Dudley, (2002).
Theorem 1.
Let Assumption 1 hold. Partition $W=(Z^{\prime},1)^{\prime}\in\mathbb{R}^{K}$. Suppose that the support of $Z$ is convex and has non-empty interior. Let Assumption 2 hold.
Then $b=\mu\beta$ for some $\mu>0$ if and only if $R(b)=0$, where
$R(b)=P\left({\rm sgn}\left(E(Y_{1}-Y_{0}\mid X)\right)\neq{\rm sgn}(W^{\prime}b)\right).$ (3)
Therefore, $\beta$ is identified up to scale.
The requirement on the distribution of $Z$ is mild. Let $\mathcal{Z}$ be the support of $Z$. First, we do not require $\mathcal{Z}$ to be an unbounded set. Second, the distribution of $Z$ does not have to admit a density and “atoms” (or point masses) are allowed. Theorem 1 imposes this via the convexity and non-empty interior of $\mathcal{Z}$. When $\mathcal{Z}$ is not convex, the result is still useful: as long as $\mathcal{Z}$ contains a convex subset with non-empty interior, we can apply the result on this subset by restricting the sample, i.e., the sign saturation condition holds on the restricted sample. If there are multiple discrete variables, then we set $Z$ to be the difference in continuous variables, see Section 2.2.
By Theorem 1, to check whether $b$ is a rescaled version of the true $\beta$, we only need to check $R(b)$ from the distribution of the observed data. Notice that the identification does not assume that $u_{t}$’s are independent across $t$; this is because Assumption 1 only requires them to have the same marginal distribution given $(X,\alpha)$. To give some intuition on the identifying power of sign saturation under bounded $\mathcal{Z}$, we use the following simple example to illustrate why the criterion function $R(\cdot)$ fails to deliver identification when Assumption 2 is not satisfied.
Example 1.
Suppose that $W=(Z,1)$, where $Z$ is a continuous random variable with support $[-1,1]$. For simplicity, we consider $\beta=(\beta_{1},\beta_{2})^{\prime}$ and $b=(b_{1},b_{2})^{\prime}$ with $\beta_{1},b_{1}>0$. The question is whether we can derive $b=\mu\beta$ for some $\mu>0$ from $R(b)=0$. Here, identifying $\beta$ up to scale is equivalent to identifying $\beta_{2}/\beta_{1}$. After straight-forward calculations, we can see that
$R(b)=P\left(Z\in\left(-\frac{\beta_{2}}{\beta_{1}},-\frac{b_{2}}{b_{1}}\right)\bigcup\left(-\frac{b_{2}}{b_{1}},-\frac{\beta_{2}}{\beta_{1}}\right)\right).$
Suppose that the sign saturation condition fails, say $P(W^{\prime}\beta>0)=0$, which implies that $-\beta_{2}/\beta_{1}\geq 1$. In this case, the set $\left(-\frac{\beta_{2}}{\beta_{1}},-\frac{b_{2}}{b_{1}}\right)\bigcup\left(-\frac{b_{2}}{b_{1}},-\frac{\beta_{2}}{\beta_{1}}\right)$ has empty intersection with $[-1,1]$ whenever $-b_{2}/b_{1}>1$. In other words, if $P(W^{\prime}\beta>0)=0$, then $R(b)=0$ whenever $-b_{2}/b_{1}>1$; as a result, we cannot always verify $b_{2}/b_{1}=\beta_{2}/\beta_{1}$ through $R(b)$.
On the other hand, if the sign saturation condition holds, then we have $-1<\beta_{2}/\beta_{1}<1$. In this case, the intersection of $\left(-\frac{\beta_{2}}{\beta_{1}},-\frac{b_{2}}{b_{1}}\right)\bigcup\left(-\frac{b_{2}}{b_{1}},-\frac{\beta_{2}}{\beta_{1}}\right)$ and $[-1,1]$ always has non-empty interior whenever $b_{2}/b_{1}\neq\beta_{2}/\beta_{1}$. Consequently, $R(b)=0$ if and only $b_{2}/b_{1}=\beta_{2}/\beta_{1}$. ∎
The following example demonstrates how Theorem 1 gives identification conditions for $\beta$ with interaction terms.
Example 2 (Interaction terms).
Let $D_{t}=\mathbf{1}\{t=1\}$ and consider a continuous variable $H_{t}$. Suppose that we are interested in the treatment effect of $D_{t}$ controlling for $H_{t}$. For a more flexible specification, we include the interaction $D_{t}H_{t}$. Then $X_{t}=(D_{t},H_{t},D_{t}H_{t})^{\prime}\in\mathbb{R}^{3}$ with the corresponding partition $\beta=(\beta_{1},\beta_{2},\beta_{3})^{\prime}\in\mathbb{R}^{3}$. For simplicity, suppose that the support for $(H_{0},H_{1})^{\prime}$ is $[-1,1]\times[-1,1]$. Then $Z=(H_{1},D_{1}H_{1})^{\prime}-(H_{0},D_{0}H_{0})^{\prime}=(H_{1}-H_{0},H_{1})^{\prime}$. One can easily verify that $\mathcal{Z}$ is convex and has non-empty interior. [^2] To see the convexity, simply notice that $Z=\begin{pmatrix}-1&1\\ 0&1\end{pmatrix}\begin{pmatrix}H_{0}\\ H_{1}\end{pmatrix}$ and $(H_{0},H_{1})^{\prime}$ has a convex support. By Lemma 1, straight-forward calculations show that the sign saturation condition becomes $|\beta_{2}+\beta_{3}|+|\beta_{3}|>|\beta_{1}|$. For example, this condition holds when $\beta_{1}=\beta_{3}=0$ and $\beta_{2}\neq 0$. Therefore, to test the null hypothesis of “zero effects whatsoever”, it suffices to have a relevant control variable. ∎
We now discuss an important special case with only continuous covariates.
Corollary 1.
Let Assumption 1 hold. If the interior of the support of $X_{1}-X_{0}$ contains $\boldsymbol{0}_{\dim(X_{t})}$ (zero in $\mathbb{R}^{\dim(X_{t})})$ and $\beta\neq\boldsymbol{0}_{\dim(X_{t})}$, then $\beta$ is identified up to scale.
A comparison between Example 1 and Corollary 1 highlights the role of discrete variables in identification. If all the covariates are continuous and the change in covariates contains zero in the support, then we can generally expect identification (up to scale) of $\beta$, regardless of whether the covariates are bounded. In contrast, if the covariates include discrete variables (such as the time fixed effect in Chamberlain, (2010)), we no longer have such a generic guarantee of identification.
2.2 Extension to multiple discrete regressors
In this subsection, we consider the case of multiple discrete regressors. [^3] I thank Xavier D’Haultfœuille for a discussion of this case. We partition $X_{t}=(X_{(1),t}^{\prime},X_{(2),t}^{\prime})^{\prime}$, where $X_{(1),t}\in\mathbb{R}^{K_{1}}$ are discrete regressors and $X_{(2),t}\in\mathbb{R}^{K_{2}}$ are continuous regressors. We can partition the corresponding $W=X_{1}-X_{0}=(D^{\prime},Z^{\prime})^{\prime}$ with $D=X_{(1),1}-X_{(1),0}$ and $Z=X_{(2),1}-X_{(2),0}$. Let $\mathcal{D}$ denote the support of $D$. We show that identification still holds under a conditional version of sign saturation.
Theorem 2.
Let Assumption 1 hold. Suppose that there exist linearly independent $d_{1},...,d_{K_{1}}\in\mathcal{D}$ with such that for $j\in\{1,...,K_{1}\}$,
(1) both $P(E(Y_{1}-Y_{0}\mid X,D=d_{j})>0)$ and $P(E(Y_{1}-Y_{0}\mid X,D=d_{j})<0)$ are strictly positive
(2) the support of $Z$ conditional on $D=d_{j}$ is convex and has non-empty interior.
Then $\beta$ is identified up to scale.
We note that $|\mathcal{D}|$ can be much larger than $K_{1}$. For example, if $X_{(1),t}$ represents $K_{1}$ binary variables, then the support of $X_{(1),t}$ is $\{0,1\}^{K_{1}}$ and $\mathcal{D}=\{-1,0,1\}^{K_{1}}$, which means $|\mathcal{D}|=3^{K_{1}}$. Theorem 2 is convenient in that it does not require the sign saturation to hold conditional on $D=d$ for every $d\in\mathcal{D}$. It suffices to require this for $K_{1}$ linearly independent elements of $\mathcal{D}$. If $K_{1}=1$ ($X_{(1),t}$ is a time dummy), then this requirement reduces to Assumption 2. Theorem 2 also allows for more complicated cases as demonstrated in the following example.
Example 3 (Time dummy and categorical variables).
Suppose that $R_{t}$ is a categorical variable taking values in $\{1,...,K_{1}\}$. Consider $X_{(1),t}=(\mathbf{1}\{t=1\},\mathbf{1}\{R_{t}=1\},...,\mathbf{1}\{R_{t}=K_{1}-1\})^{\prime}\in\mathbb{R}^{K_{1}}$, i.e., a time dummy and $K_{1}-1$ dummy variables for $R_{t}$. Then we can write the support of $X_{(1),t}\in\mathbb{R}^{K_{1}}$ as $\{\mathbf{1}\{t=1\}\}\times\{e_{1},...,e_{K_{1}-1},\boldsymbol{0}_{K_{1}-1}\}$, where $e_{j}$ is the $j$-th column of the $(K_{1}-1)\times(K_{1}-1)$ identity matrix and $\boldsymbol{0}_{K_{1}-1}=(0,...,0)^{\prime}\in\mathbb{R}^{K_{1}-1}$. Let us assume that the support of $(R_{0},R_{1})^{\prime}$ contains $(K_{1},j)$ for any $j\in\{1,...,K_{1}\}$. Then $\mathcal{D}$ contains $\{1\}\times\{e_{1},...,e_{K_{1}-1},\boldsymbol{0}_{K_{1}-1}\}$. In other words, $\mathcal{D}$ contains all the columns of the following matrix:
$\begin{pmatrix}1&1&1&\cdots&1\\ 0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{pmatrix}\in\mathbb{R}^{K_{1}\times K_{1}}.$
We can easily see that the determinant of this upper triangular matrix is equal to one. Therefore, the above matrix has full rank, which means that $\mathcal{D}$ contains $K_{1}$ linearly independent elements. Then $\beta$ is identified up to scale if for every $j\in\{1,...,K_{1}\}$, $P(E(Y_{1}-Y_{0}\mid Z,R_{0}=K_{1},R_{1}=j)>0)$ and $P(E(Y_{1}-Y_{0}\mid Z,R_{0}=K_{1},R_{1}=j)<0)$ are strictly positive and the support of $Z$ conditional on $(R_{0},R_{1})=(K_{1},j)$ is convex and has non-empty interior.∎
2.3 Is the sign saturation condition is also necessary?
It turns out that in the many situations, sign saturation characterizes identifiability. To illustrate this point, we consider the setting studied by Chamberlain, (2010) and assume that $u_{t}$ is independent of $(X,\alpha)$ and is from a known distribution. [^4] To establish necessity, it is enough to consider the case in which $u_{t}$ is independent of $(X,\alpha)$. This is similar to showing minimax lower bounds. If identification cannot be guaranteed even with the additional assumption of independence between $u$ and $(X,\alpha)$, then it is definitely not guaranteed without this assumption. We show that without sign saturation, identification fails at every point unless the distribution of $u_{t}$ is from a special class.
Suppose that in time period $t\in\{0,1\}$, we observe $(Y_{t},X_{t})$, where $X_{t}=(X_{t,1}^{\prime},X_{t,2})^{\prime}$ with $X_{t,1}\in\mathbb{R}^{K-1}$ and $X_{t,2}=\mathbf{1}\{t=1\}$. Suppose that $u_{t}$ is independent of $(X,\alpha)$ and has c.d.f $F(\cdot)$. Then from the data, the distribution of $Y$ given $(X,\alpha)$ is determined by the vector
$L(X;\beta,\alpha)$ $=\begin{pmatrix}F(X_{0,1}^{\prime}\beta_{1}+\alpha)\\ F(X_{1,1}^{\prime}\beta_{1}+\beta_{2}+\alpha)\\ F(X_{0,1}^{\prime}\beta_{1}+\alpha)\cdot F(X_{1,1}^{\prime}\beta_{1}+\beta_{2}+\alpha)\end{pmatrix}$
$=\begin{pmatrix}F(X_{0,1}^{\prime}\beta_{1}+\alpha)\\ F(X_{0,1}^{\prime}\beta_{1}+\alpha+W^{\prime}\beta)\\ F(X_{0,1}^{\prime}\beta_{1}+\alpha)\cdot F(X_{0,1}^{\prime}\beta_{1}+\alpha+W^{\prime}\beta)\end{pmatrix},$ (4)
where $\beta=(\beta_{1}^{\prime},\beta_{2})$ is partitioned as $\beta_{1}\in\mathbb{R}^{K-1}$ and $\beta_{2}\in\mathbb{R}$, and $W:=X_{1}-X_{0}=(Z^{\prime},1)^{\prime}$ with $Z=X_{1,1}-X_{0,1}$. As pointed out in Chamberlain, (2010), here we should aim for identification of $\beta$, not just up to scaling, because “our scale normalization is built in to the given specification for the $u_{t}$ distribution”. Let $\mathcal{Z}$ be the support of $Z$, i.e., again the smallest closed set with the full probability mass. Since $W=(Z^{\prime},1)^{\prime}$, the support of $W$ is $\mathcal{W}=\mathcal{Z}\times\{1\}$.
The special class of functions is best described in terms of a transformation of $F(\cdot)$. Define the function $G(\cdot)=\ln\frac{F(\cdot)}{1-F(\cdot)}$. Let $\dot{G}(\cdot)$ denote the derivative of $G(\cdot)$. If $F(\cdot)$ is the logistic function, then $G(\cdot)$ is an affine function and $\dot{G}(\cdot)$ is a constant function. It turns out that if we are willing to rule out cases with $\dot{G}(\cdot)$ being a periodic function, [^5] A function $h(\cdot)$ on $\mathbb{R}$ is a periodic function if there exists $c\neq 0$ such that $h(c+t)=h(t)$ for any $t\in\mathbb{R}$. In this case, $c$ is said to be a period of $h(\cdot)$. Notice that we assume that a period has to be non-zero; otherwise, every function would be a periodic function with a period being zero. then sign saturation is a necessary condition for identification. This special class includes the logistic function, whose $\dot{G}(\cdot)$ is a constant function and is clearly periodic. (Any real number is a period of a constant function.)
[…]
Definition 1 (Identification failure at $\beta$).
[…]
Notice that $X_{0,1}^{\prime}\beta_{1}$ can be absorbed by $\alpha$ since the distribution of $\alpha$ is allowed to have arbitrary dependence on $x$. Without loss of generality, we view the entire $X_{0,1}^{\prime}\beta_{1}+\alpha$ as the fixed effects to simplify notations. Then we can restate Definition 1 as follows.
Definition 2.
We say that the identification fails at $\beta$ if there exists $b\neq\beta$ such that $\mathcal{A}(w^{\prime}\beta)\bigcap\mathcal{A}(w^{\prime}b)\neq\emptyset$ for any $w\in\mathcal{W}$, where for any $t\in\mathbb{R}$, $\mathcal{A}(t)={\rm conv}\{p(t,\alpha):\alpha\in\mathbb{R}\}$ and
$p(t,\alpha):=\begin{pmatrix}F(\alpha)\\ F(t+\alpha)\\ F(\alpha)\cdot F(t+\alpha)\end{pmatrix}.$
Perhaps the simplest way to see the equivalence between the two definitions is to notice that ${\rm conv}\{L(x;\beta,\alpha):\alpha\in\mathbb{R}\}=\mathcal{A}(w^{\prime}\beta)$, where $w=x_{1}-x_{0}$. We now define the parameter space in which the sign saturation fails: +=$\mathcal{B}_{+}=\{\beta\in\mathbb{R}^{K}:\ w^{\prime}\beta>0\ \forall w\in\mathcal{W}\}$. The main result for the necessity is the following.
Theorem 3.
Suppose that $\mathcal{Z}$ is bounded. Suppose that $F(\cdot)$ is continuously differentiable with support on $\mathbb{R}$. If $\dot{G}(\cdot)$ is not a periodic function, then the identification fails at every point in $\mathcal{B}_{+}$.
We compare this with Theorem 1 of Chamberlain, (2010), which states that if $F(\cdot)$ is outside a special class, identification fails in an open neighborhood. Theorem 3 complements this result by showing that if $F(\cdot)$ is outside a special class, identification fails at every point, not just in an open neighborhood. This is also why the proof of Theorem 3 follows a very different strategy from the proof of Theorem 1 of Chamberlain, (2010); the latter essentially only needs to find one point at which the identification fails whereas we need to show that identification fails universally in $\mathcal{B}_{+}$.
The special class in Chamberlain, (2010) only includes logistic functions, but here our special class is larger and includes all functions with periodic $\dot{G}(\cdot)$. This enlargement of the special class cannot be avoided because we show that there are instances of non-logistic $F(\cdot)$ for which identification holds, e.g., $F(t)=[1+\exp(-2t-\sin(t))]^{-1}$. We can push our arguments further and show that the identification under periodic $\dot{G}(\cdot)$ is “robust” and is based on a set with strictly positive probability mass.
Theorem 4.
Suppose that $\dot{G}(\cdot)$ is a continuous periodic function that is non-constant. Let $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}\in\mathcal{B}_{+}$. Then
(1) there exists a minimal positive period $\eta>0$ for $\dot{G}(\cdot)$, i.e., the set $\{a>0:\ \dot{G}(a+t)=\dot{G}(t)\ \forall t\in\mathbb{R}\}$ has a smallest element.
(2) If $(z^{\prime}\beta_{1}+\beta_{2})/\eta$ is an integer for some $z$ in the interior of $\mathcal{Z}$, then $\beta$ is identified; moreover, for any $b\in\mathcal{B}_{+}$ with $b\neq\beta$, $P\left(\mathcal{A}(W^{\prime}\beta)\bigcap\mathcal{A}(W^{\prime}b)=\emptyset\right)>0$, where $W=(Z^{\prime},1)^{\prime}$.
(3) If $\mathcal{Z}$ is compact and has non-empty interior, then there exists an open ball $D\subset\mathcal{B}_{+}$ such that every point in $D$ is identified.
By Theorem 4, as long as $\dot{G}(\cdot)$ is periodic and $z^{\prime}\beta_{1}+\beta_{2}$ is in $\eta\cdot\mathbb{N}$ for some interior point $z$ ($\mathbb{N}$ denotes the set of positive integers), we still have identification. This complements Theorem 1 of Chamberlain, (2010) in an interesting way. For non-logistic $F(\cdot)$, it is true that the identification fails at every point in an open set. On the other hand, we show that the identification also holds at every point in another open set for periodic $\dot{G}(\cdot)$. It is worth noting that in Theorem 4, the identification of $\beta$ is robust in the sense that it is not based on a small number of “unimportant points” in $\mathcal{W}$; the points in $\mathcal{W}$ that allow us to identify $\beta$ have strictly positive probability mass. Therefore, the truly hopeless cases for identification are those with non-periodic $\dot{G}(\cdot)$. Finally, for the case of compact and convex $\mathcal{Z}$ with non-empty interior, we summarize our results and the existing literature in Table 1.
Table 1: Identification and sign saturation
For bounded and convex $\mathcal{Z}$ with non-empty interior:
Sign saturation holds Sign saturation does not hold
$\dot{G}(\cdot)$ non-periodic
ID holds at every point
(Theorem 1)
ID fails at every point
(Theorem 3)
$\dot{G}(\cdot)$ periodic
but non-constant
ID holds at every point
(Theorem 1)
ID holds in an open set
(Theorem 4)
but fails in another open set
(Chamberlain, (2010))
$\dot{G}(\cdot)$ constant
(logistic $F$)
ID holds at every point
(Chamberlain, (1980)
Davezies et al., (2022))
ID holds at every point
(Chamberlain, (1980)
Davezies et al., (2022))
Exploiting the periodicity of $\dot{G}(\cdot)$ can be seen as a generalization of the conditional maximum likelihood estimator for logistic distributions. To see this, we notice that
$\frac{P(Y_{1}=1,Y_{0}=0\mid X,\alpha)}{P(Y_{1}=0,Y_{1}=1\mid X,\alpha)}=\exp\left[G(W^{\prime}\beta+\alpha)-G(\alpha)\right].$
Under logistic errors, $\dot{G}(\cdot)$ is a constant and thus $G(W^{\prime}\beta+\alpha)-G(\alpha)$ does not depend on $\alpha$ for any $W$. If $\dot{G}(\cdot)$ is a periodic function with the smallest positive period $\eta>0$, then $G(W^{\prime}\beta+\alpha)-G(\alpha)$ also does not depend on $\alpha$ whenever $W^{\prime}\beta/\eta$ is an integer. Since $W=(Z^{\prime},1)^{\prime}$, we only need to have an interior point $z$ such that $(z^{\prime},1)^{\prime}\beta/\eta$ is an integer.
Remark 2.
As we have seen, the sign saturation condition guarantees the identification of $\beta$ up to scale. Does it also guarantee the parametric rate for estimation? The answer is no because Theorem 2 of Chamberlain, (2010) shows that the information bound is always zero outside the logistic case, regardless of the identification status. This highlights a key difference between identification and the parametric rate in estimation. The question of identification depends on the sign saturation condition, whereas the root-$n$ estimability depends on the semiparametric efficiency calculation or the functional projection argument developed by Bonhomme, (2012).
3 Implications on marginal effects
The identification of $\beta$ is directly related to the identification of marginal effects in panel models with binary responses. The bounds on the marginal effects often assume point identification of $\beta$, see e.g., Davezies et al., (2022), Liu et al., (2024) and Theorem 6 of Chernozhukov et al., (2013).
Here, we explore an important link between $\beta$ and the treatment effects through the sign of components of $\beta$. Let us consider the setting of Section 2.3: (1) the errors $u_{t}$’s are i.i.d across $t$ with distribution $F(\cdot)$ and are independent of $(X,\alpha)$ and (2) $X_{2,t}$ is binary with $X_{2,t}=\mathbf{1}\{t=1\}$. We can interpret $X_{2,t}$ as a binary treatment; in period $t=0$, no one is treated and in period $t=1$, everyone is treated. For $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}$, the sign of $\beta_{2}$ corresponds to the sign of common measures of treatment effects. For example, the average partial effect effect at $X_{1,t}=z$ is
$\Delta_{APE}(z)=P(z^{\prime}\beta_{1}+\beta_{2}+\alpha\geq u_{t})-P(z^{\prime}\beta_{1}+\alpha\geq u_{t}),$
see e.g., Chernozhukov et al., (2013).
Although the magnitude of average partial effect is often not point identified (see e.g., Honoré and Tamer, (2006) and Chernozhukov et al., (2013)), there is hope that the sign of the effect is point identified, which means that the identified set contains only positive numbers, only negative numbers or only zero. Under Assumption 1, the support of $u_{t}$ is $\mathbb{R}$, which means that ${\rm sgn}(\Delta_{APE}(z))={\rm sgn}(\beta_{2})$ for any $z$. Hence, identifying ${\rm sgn}(\Delta_{APE}(z))$ is equivalent to identifying ${\rm sgn}(\beta_{2})$. In the literature, there are also terms such as average treatment effects or ceteris paribus effects, e.g., Hoderlein and White, (2012) and Chernozhukov et al., (2015). Under the assumptions of a linear index, the sign of these other measures of treatment effects is also the sign of $\beta_{2}$.
Remark 3.
There are alternative ways of defining treatment effects but due to the single-index nature of the model, the sign of other notions of treatment effects is still related to ${\rm sgn}(\beta_{2})$. For example, consider the individual effect $\mathbf{1}\{X_{1,t}\beta_{1}+\beta_{2}+\alpha\geq u_{t}\}-\mathbf{1}\{X_{1,t}\beta_{1}+\alpha\geq u_{t}\}$. Clearly, this effect is negative with zero probability if and only if $\beta_{2}>0$.
We now show that the sign saturation condition is necessary for this purpose. To formally state the result, we rephrase Definition 1.
Definition 3.
[…]
Recall that +=$\mathcal{B}_{+}=\{\beta\in\mathbb{R}^{K}:\ w^{\prime}\beta>0\ \forall w\in\mathcal{W}\}$ is the set of parameter values that do not satisfy the sign saturation condition ($E(Y_{1}-Y_{0}\mid X)$ always positive). Consider two subsets $\mathcal{B}_{+,0}=\{\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}\in\mathcal{B}_{+}:\ \beta_{2}=0\}$ and +,+=$\mathcal{B}_{+,+}=\{\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}\in\mathcal{B}_{+}:\ \beta_{2}>0\}$, which denote the set of parameter values that imply zero effects and positive effects of $X_{2,t}$, respectively. The following result states the identification failure of ${\rm sgn}(\beta_{2})$.
Theorem 5.
Suppose that $\mathcal{Z}$ is bounded. Suppose that $F(\cdot)$ is continuously differentiable such that $\dot{G}(\cdot)$ is not a periodic function. Then every point in $\mathcal{B}_{+,0}$ is observationally equivalent to a point in $\mathcal{B}_{+,+}$.
By Theorem 1, the sign saturation condition guarantees point identification of $\beta$ up to scale and thus point identification of ${\rm sgn}(\beta_{2})$. By Theorem 5, when the sign saturation condition fails, we cannot distinguish between $\Delta_{APE}(z)=0$ and $\Delta_{APE}(z)>0$ unless $\dot{G}(\cdot)$ is periodic. Therefore, unless $\dot{G}(\cdot)$ is periodic, the sign saturation condition is sufficient and necessary to guarantee point identification of the sign of the marginal effects of $X_{2,t}$.
4 Assessing the sign saturation condition
We have seen that the sign saturation condition is sufficient and necessary for the identification. The conditional mean function $\phi(X)=E(Y_{1}-Y_{0}\mid X)$ is identified. Therefore, ideally the sign saturation condition can be checked in data. Here, we provide a simple check that does not involve nonparametric estimation of $\phi$.
By Lemma 1, $P(\phi(X)\leq 0)=1$ if and only if $P(W^{\prime}\beta\leq 0)=1$. Hence, we define $\rho(q)=E\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})$ for $q\in\mathbb{R}^{K}$. Here, we notice that we can replace $\mathbb{R}^{K}$ with $[-1,1]^{K}$ or any set that contains an open neighborhood of zero. It turns out that the sign saturation condition can be written in terms of $\rho(\cdot)$ once we observe
$\rho(\beta)=E\left(\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)\right)\quad{\rm and}\quad\rho(-\beta)=E\left(\mathbf{1}\{\phi(X)\leq 0\}\cdot\phi(X)\right).$
We give the formal statement below.
Lemma 2.
Let Assumption 1 hold. Then
$\sup_{q\in\mathbb{R}^{K}}\rho(q)=E\max\left\{\phi(X),0\right\}\ {\rm and}\ \inf_{q\in\mathbb{R}^{K}}\rho(q)=E\min\left\{\phi(X),0\right\}.$
Define $\tau_{*}=\min\{\tau_{1},-\tau_{2}\}$ with $\tau_{1}=\sup_{q\in\mathbb{R}^{K}}\rho(q)$ and $\tau_{2}=\inf_{q\in\mathbb{R}^{K}}\rho(q)$. By Lemma 2, measuring $\tau_{*}$ can give us some indication of whether (or how “well”) the sign saturation condition holds. In particular, the sign saturation fails if and only if $\tau_{*}=0$. From the data, we can construct a one-sided confidence interval for $\tau_{*}$. The natural estimate for $\tau_{*}$ is $\hat{\tau}_{*}=\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}$, where $\hat{\tau}_{1}=\sup_{q\in\mathbb{R}^{K}}\hat{\rho}_{n}(q)$, $\hat{\tau}_{2}=\inf_{q\in\mathbb{R}^{K}}\hat{\rho}_{n}(q)$ and
$\hat{\rho}_{n}(q)=n^{-1}\sum_{i=1}^{n}\mathbf{1}\{W_{i}^{\prime}q\geq 0\}(Y_{i,1}-Y_{i,0}).$
In the proof of Theorem 6 below, we show that
$\sqrt{n}\left(\hat{\tau}_{*}-\tau_{*}\right)\leq\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)\right)\cdot\mathbf{1}\{\tau_{1}\leq-\tau_{2}\}+\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))\right)\cdot\mathbf{1}\{\tau_{1}>-\tau_{2}\},$
where $S_{n}(q)=\sqrt{n}(\hat{\rho}_{n}(q)-\rho(q))$. By the standard arguments of empirical processes, $S_{n}$ converges to a mean-zero Gaussian process, which means that $\sup_{q\in\mathbb{R}^{K}}S_{n}(q)$ and $\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))$ have the same distribution. Thus, a simple $(1-\alpha)$-confidence interval for $\tau_{*}$ can be obtained once we approximate the $(1-\alpha)$ quantile of $\sup_{q\in\mathbb{R}^{K}}S_{n}(q)$. This motivates the following bootstrapping algorithm.
Algorithm 1.
For a $(1-\alpha)$-confidence interval for $\tau_{*}$:
  1. 1.
    Collect data $\{(W_{i},Y_{i,1},Y_{i,0})\}_{i=1}^{n}$.
  2. 2.
    Compute the estimate $\hat{\tau}_{*}=\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}$.
  3. 3.
    Draw a random sample $\{(\tilde{W}_{i},\tilde{Y}_{i,1},\tilde{Y}_{i,0})\}_{i=1}^{n}$ with replacement from the data and compute $\sup_{q\in\mathbb{R}^{K}}\tilde{S}_{n}(q)$, where $\tilde{S}_{n}(q)=\sqrt{n}\left(n^{-1}\sum_{i=1}^{n}\mathbf{1}\{\tilde{W}_{i}^{\prime}q>0\}(\tilde{Y}_{i,1}-\tilde{Y}_{i,0})-\hat{\rho}_{n}(q)\right)$.
  4. 4.
    Repeat the previous step many times and compute $c_{1-\alpha}$, the $(1-\alpha)$ quantile of $\sup_{q\in\mathbb{R}^{K}}\tilde{S}_{n}(q)$. [^6] In practice, optimization over $q\in\mathbb{R}^{K}$ can be done by a grid search over $N$ points on the unit sphere $\{v\in\mathbb{R}^{K}:\ \|v\|_{2}=1\}$. The computation seems reasonably fast for a normal sample size. For example, in the setting of Section 5 with $\beta_{2}=1$, $n=5000$ and grid size $N=2000$, computing $\sup_{q}\tilde{S}_{n}(q)$ with 500 bootstrap samples take less than 17 seconds in Matlab on a 2022 MacBook Pro laptop. Further speedups are possible depending on $n$, $N$ and available memory.
  5. 5.
[…]
Notice that computing $\hat{\tau}_{1}$ is equivalent to computing the maximum score estimator [^7] Notice that maximizing $\hat{\rho}_{n}(q)$ is equivalent to maximizing $n^{-1}\sum_{i=1}^{n}{\rm sgn}(W_{i}^{\prime}q)(Y_{i,1}-Y_{i,0})$ because ${\rm sgn}(W_{i}^{\prime}q)=-1+2\cdot\mathbf{1}\{W_{i}^{\prime}q\geq 0\}$. and all the existing computational algorithms and software for the maximum score estimator can be used. Finding $\hat{\tau}_{2}$ also reduces to computing the maximum score estimator once we swap $Y_{1}$ and $Y_{0}$. The above algorithm can be justified by the following result.
Theorem 6.
Let Assumption 1 hold. Assume that $P(Y_{1}=Y_{0})<1$. Then
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}(\hat{\tau}_{*}-\tau_{*})>c_{1-\alpha}\right)\leq\alpha.$
Note that the cube-root asymptotics typically associated with the maximum score estimator (see e.g., Kim and Pollard, (1990) and Seo and Otsu, (2018)) does not arise here. The reason is that $\hat{\tau}_{1}$ is the maximum of $\hat{\rho}_{n}(\cdot)$, rather than the argmax. The cube-root asymptotics of the maximum score estimator is due to the non-standard rate of some terms in the expansion for analyzing the argmax. Fortunately, we do not have to deal with such terms for our purpose.
It is worth pointing out that Algorithm 1 is asymptotically exact in the sense that there exists a data-generating process under which the asymptotic coverage probability is exactly $\alpha$. For example, if $\rho(q)=0$ for any $q$, then $\sqrt{n}(\hat{\tau}_{*}-\tau_{*})=\sup_{q\in\mathbb{R}^{K}}S_{n}(q)$ and $P\left(\sqrt{n}(\hat{\tau}_{*}-\tau_{*})>c_{1-\alpha}\right)\rightarrow\alpha$.
Our results characterize what drives the identification and introduce $\tau_{*}$ as a measure for the identification strength. This theoretical insight motivates a preliminary diagnostic step in applied work. Reporting the confidence interval of $\tau_{*}$ in Algorithm 1 can offer important insight on the identification strength. For example, if this confidence interval contains zero, then one should be concerned about identification failure.
However, testing $\tau_{*}=0$ might not detect all the problematic cases. In practice, the identification might not be clear-cut and additional caution is warranted if we worry about the “weak” identification scenario, which can complicate econometric analysis due to the generic issue of post-selection inference, see e.g., Leeb and Pötscher, (2005).
Fortunately, since $\tau_{*}$ can be learned from the data, one option is to proceed with the identification-based inference (i.e., cube-root asymptotics) only when the identification is strong enough. For instance, one might adopt a robust default inference method and switch to the cube-root approach only when a test fails to reject $H_{0}:\ \tau_{*}\geq\tau_{0}$, where $\tau_{0}>0$ is a pre-specified fixed threshold. This is uniformly valid asymptotically as it rules out $\tau_{*}$ being “local” to zero.
This strategy is conceptually analogous to a common treatment of weak instruments in the linear instrumental variable (IV) models: low correlation between the instrument and the exogenous variable suggests potential identification failure and one solution is to proceed with the classical asymptotics only when this correlation is large enough, e.g., measured by the first-stage $F$-statistic (Stock and Yogo,, 2002).
This raises a natural question: what should this default identification-robust method be? In the IV literature, approaches such as the Anderson-Rubin test have been developed. In our model, one could consider inverting the maximum score criterion function, in a manner similar to the Anderson–Rubin approach. While more refined methods may be possible, they would require substantial additional analysis on asymptotic theory, which lies beyond the current focus on characterizing identification and are left for future research.
5 Numerical illustration
We have seen that $\tau_{*}$ is a measure of the sign saturation condition and thus serves a gauge for identification. Using Monte Carlo simulations, we now illustrate how $\tau_{*}$ relates to the identification strength and the accuracy of the cube-root asymptotics.
Consider the model in (1) with $X_{t}=(\mathbf{1}\{t=0\},X_{2,t})^{\prime}$ for $t\in\{0,1\}$, where $X_{2,t}$ is from the uniform distribution on $[-1,1]$, $u_{t}$ is from the standard normal distribution and $\alpha_{t}=(X_{2,0}+X_{2,1})/2$. Here, $X_{2,0}$, $X_{2,1}$, $u_{0}$ and $u_{1}$ are mutually independent. We set $\beta=(1,\beta_{2})^{\prime}$. Therefore, $W^{\prime}\beta=1+\beta_{2}(X_{2,1}-X_{2,0})$. Clearly, the support of $W^{\prime}\beta$ is $[-2|\beta_{2}|+1,2|\beta_{2}|+1]$, which means that the sign saturation holds if and only if $|\beta_{2}|>0.5$. In Figure 3, we plot $\tau_{*}$ (computed using $\hat{\tau}_{*}$ with $n=10^{8}$) as a function of $\beta_{2}$. It confirms the intuition that $\beta_{2}=0.5$ corresponds to $\tau_{*}=0$, which means that the sign saturation condition fails. Higher value of $\beta_{2}$ corresponds to a higher degree to which the sign saturation condition holds.
Figure 3: Coefficient and identification
[…]
Figure 4: Distribution of $n^{1/3}(\hat{\beta}_{2}-\beta_{2})$ for $\beta_{2}=0.6$ (left) and $\beta_{2}=1$ (right)
Appendix A Appendix: proofs of theoretical results
A.1 Proof of Theorem 1
We define the following notations that will be used throughout this subsection (proof of Theorem 1). For any set $A$, $\overline{A}$ denotes the closure of $A$ and $A^{\circ}$ denotes the interior of $A$. Let $\mathcal{Z}$ be the support of $Z$. We partition $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}$ and $b=(b_{1}^{\prime},b_{2})^{\prime}$ with $\beta_{1},b_{1}\in\mathbb{R}^{K-1}$ and $\beta_{2},b_{2}\in\mathbb{R}$. Then $W^{\prime}\beta=Z^{\prime}\beta_{1}+\beta_{2}$. Define the following sets: $A_{1}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}\beta_{1}+\beta_{2}>0\}$, $A_{2}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}b_{1}+b_{2}<0\}$, $B_{1}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}\beta_{1}+\beta_{2}<0\}$ and $B_{2}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}b_{1}+b_{2}>0\}$. Before we prove Theorem 1, we establish some auxiliary results.
Lemma 3.
Suppose that $A$ is an open set (in $\mathbb{R}^{K-1}$) such that $A\bigcap\mathcal{Z}\neq\emptyset$. Then $P(Z\in A)>0$.
Proof.
Recall that $\mathcal{Z}$ is the smallest closed set such that $P(Z\in\mathcal{Z})=1$, see the definition of the support of a measure on page 227 of Dudley, (2002).
Suppose that $P(Z\in A)=0$. Then $P(Z\in\mathcal{Z}\backslash A)=1$. Since $A$ is open, $\mathcal{Z}\backslash A$ is a closed set. Notice that $\mathcal{Z}\backslash A$ is a proper subset of $\mathcal{Z}$ because $A\bigcap\mathcal{Z}$ is non-empty. Then $\mathcal{Z}\backslash A$ is a closed set with full probability mass and is a smaller set than $\mathcal{Z}$, contradicting the assumption that $\mathcal{Z}$ is the support of $Z$. ∎
Lemma 4.
Suppose that $B$ is an open set (in $\mathbb{R}^{K-1}$) such that $\overline{B}\bigcap\mathcal{Z}^{\circ}\neq\emptyset$. Then $P(Z\in B)>0$.
Proof.
Let $A=B\bigcap\mathcal{Z}^{\circ}$. By the openness of $B$ and $\mathcal{Z}^{\circ}$, $A$ is open. The desired result follows by Lemma 3 once we show that $A$ is non-empty.
We do so by contradiction. Suppose $B\bigcap\mathcal{Z}^{\circ}=\emptyset$. Then $B\subseteq(\mathcal{Z}^{\circ})^{c}$. Since $\mathcal{Z}^{\circ}$ is open, $(\mathcal{Z}^{\circ})^{c}$ is closed. Thus, $(\mathcal{Z}^{\circ})^{c}$ is a closed set containing $B$. Recall that $\overline{B}$ is the smallest closed set containing $B$. It follows that $\overline{B}\subseteq(\mathcal{Z}^{\circ})^{c}$, which means that $\overline{B}\bigcap\mathcal{Z}^{\circ}=\emptyset$. This contradicts the assumption of $\overline{B}\bigcap\mathcal{Z}^{\circ}\neq\emptyset$. Hence, $A=B\bigcap\mathcal{Z}^{\circ}$ is non-empty. ∎
Lemma 5.
Suppose that $b\neq\mu\beta$ for any $\mu>0$. If $P(Z\in A_{1})$, $P(Z\in A_{2})$, $P(Z\in B_{1})$ and $P(Z\in B_{2})$ are all strictly positive, then $\overline{A_{1}\bigcap A_{2}}=\overline{A}_{1}\bigcap\overline{A}_{2}$ and $\overline{B_{1}\bigcap B_{2}}=\overline{B}_{1}\bigcap\overline{B}_{2}$.
Proof.
We prove the first claim; the second claims follows by an analogous argument.
Since $P(Z\in A_{1})$ and $P(Z\in B_{1})$ are both strictly positive, we have that $\beta_{1}\neq 0$. To see this, observe that if $\beta_{1}=0$, then the two events $\{Z\in A_{1}\}=\{\beta_{2}>0\}$ and $\{Z\in B_{1}\}=\{\beta_{2}<0\}$ cannot both have strictly positive probability. Similarly, $P(Z\in A_{2})>0$ and $P(Z\in B_{2})$ yield yield $b_{1}\neq 0$.
First observe that $\overline{A_{1}\bigcap A_{2}}\subseteq\overline{A}_{1}\bigcap\overline{A}_{2}$ (because $A_{1}\bigcap A_{2}\subseteq\overline{A}_{1}\bigcap\overline{A}_{2}$ and $\overline{A}_{1}\bigcap\overline{A}_{2}$ is closed).
To show the other direction, we fix an arbitrary $z_{*}\in\overline{A}_{1}\bigcap\overline{A}_{2}$. We need to show that $z_{*}\in\overline{A_{1}\bigcap A_{2}}$.
Notice that $z_{*}^{\prime}\beta_{1}+\beta_{2}\geq 0$ (due to $z_{*}\in\overline{A}_{1})$ and $z_{*}^{\prime}b_{1}+b_{2}\leq 0$ (due to $z_{*}\in\overline{A}_{2})$. If $z_{*}^{\prime}\beta_{1}+\beta_{2}>0$ and $z_{*}^{\prime}b_{1}+b_{2}<0$, then $z_{*}\in A_{1}\bigcap A_{2}$, which means that $z_{*}\in\overline{A_{1}\bigcap A_{2}}$. Therefore, we only need to show that $z_{*}\in\overline{A_{1}\bigcap A_{2}}$ in the following three cases:
  • (i)
    $z_{*}^{\prime}\beta_{1}+\beta_{2}=0$ and $z_{*}^{\prime}b_{1}+b_{2}<0$
  • (ii)
    $z_{*}^{\prime}\beta_{1}+\beta_{2}>0$ and $z_{*}^{\prime}b_{1}+b_{2}=0$
  • (iii)
    $z_{*}^{\prime}\beta_{1}+\beta_{2}=0$ and $z_{*}^{\prime}b_{1}+b_{2}=0$.
Consider Case (i). Let $z_{j}=z_{*}+cj^{-1}\beta_{1}$ for $j\in\{1,2,...\}$. Since $z_{*}^{\prime}b_{1}+b_{2}<0$, we can choose a constant $c>0$ to be small enough such that $cb_{1}^{\prime}\beta_{1}<|z_{*}^{\prime}b_{1}+b_{2}|/2$, which would imply that $z_{j}^{\prime}b_{1}+b_{2}<(z_{*}^{\prime}b_{1}+b_{2})/2<0$. On the other hand, notice that $z_{j}^{\prime}\beta_{1}+\beta_{2}=(z_{*}^{\prime}\beta_{1}+\beta_{2})+cj^{-1}\|\beta_{1}\|_{2}^{2}=cj^{-1}\|\beta_{1}\|_{2}^{2}>0$ due to $\beta_{1}\neq 0$. Thus, we have constructed a sequence $\{z_{j}\}_{j=1}^{\infty}$ in $A_{1}\bigcap A_{2}$ such that $\lim_{j\rightarrow\infty}z_{j}=z_{*}$. Hence, $z_{*}$ is a limit point of $A_{1}\bigcap A_{2}$, which means that $z_{*}\in\overline{A_{1}\bigcap A_{2}}$.
Similarly, we can show that $z_{*}\in\overline{A_{1}\bigcap A_{2}}$ in Case (ii).
We now consider Case (iii). We first show that there exists $\Delta\in\mathbb{R}^{K-1}$ such that $\Delta^{\prime}\beta_{1}>0$ and $\Delta^{\prime}b_{1}<0$. We proceed by contradiction. Suppose that there does not exist $\Delta\in\mathbb{R}^{K-1}$ with $\Delta^{\prime}\beta_{1}>0$ and $\Delta^{\prime}b_{1}<0$. Then for any $\Delta\in\mathbb{R}^{K-1}$, $\Delta^{\prime}\beta_{1}>0$ implies $\Delta^{\prime}b_{1}\geq 0$. By the continuity of $\Delta\mapsto\Delta^{\prime}\beta_{1}$, $\Delta^{\prime}\beta_{1}\geq 0$ implies $\Delta^{\prime}b_{1}\geq 0$. By Farkas’s lemma (e.g., Corollary 22.3.1 of Rockafellar, (1970)), there exists $\xi\geq 0$ such that $b_{1}=\xi\beta_{1}$. Since we have proved $b_{1}\neq 0$, it follows that $\xi\neq 0$, which means that $\xi>0$. Now we have $z_{*}^{\prime}b_{1}+b_{2}=\xi z_{*}^{\prime}\beta_{1}+b_{2}=\xi(z_{*}^{\prime}\beta_{1}+b_{2}\xi^{-1})$. Since $z_{*}^{\prime}\beta_{1}+\beta_{2}=0$ and $z_{*}^{\prime}b_{1}+b_{2}=0$, it follows that $b_{2}\xi^{-1}=\beta_{2}$. Therefore, we have $b=(b_{1}^{\prime},b_{2})^{\prime}=\xi(\beta_{1}^{\prime},\beta_{2})^{\prime}$, contradicting the assumption $b\neq\mu\beta$ for any $\mu>0$. Hence, there exists $\Delta\in\mathbb{R}^{K-1}$ such that $\Delta^{\prime}\beta_{1}>0$ and $\Delta^{\prime}b_{1}<0$.
Now we choose $\Delta\in\mathbb{R}^{K-1}$ such that $\Delta^{\prime}\beta_{1}>0$ and $\Delta^{\prime}b_{1}<0$. Define $\tilde{z}_{j}=z_{*}+j^{-1}\Delta$ for $j\in\{1,2,...\}$. Then $\tilde{z}_{j}^{\prime}\beta_{1}+\beta_{2}=(z_{*}^{\prime}\beta_{1}+\beta_{2})+j^{-1}\Delta^{\prime}\beta_{1}=j^{-1}\Delta^{\prime}\beta_{1}>0$ and $\tilde{z}_{j}^{\prime}b_{1}+b_{2}=(z_{*}^{\prime}b_{1}+b_{2})+j^{-1}\Delta^{\prime}b_{1}=j^{-1}\Delta^{\prime}b_{1}<0$. Again, we have constructed a sequence $\{\tilde{z}_{j}\}_{j=1}^{\infty}$ in $A_{1}\bigcap A_{2}$ such that $\lim_{j\rightarrow\infty}\tilde{z}_{j}=z_{*}$. Hence, $z_{*}$ is a limit point of $A_{1}\bigcap A_{2}$, which means that $z_{*}\in\overline{A_{1}\bigcap A_{2}}$. ∎
Lemma 6.
Suppose that $\mathcal{Z}$ is convex and $\mathcal{Z}^{\circ}$ is non-empty. Moreover, assume that $P(Z\in A_{1})$ and $P(Z\in A_{2})$ are both strictly positive. Then at least one of $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$ and $\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}$ is not empty.
Proof.
We proceed by contradiction. Suppose that $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$ and $\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$. Clearly, $\overline{A}_{1}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}\beta_{1}+\beta_{2}\geq 0\}$ and $\overline{A}_{2}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}b_{1}+b_{2}\leq 0\}$ as well as $\overline{B}_{1}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}\beta_{1}+\beta_{2}\leq 0\}$ and $\overline{B}_{2}=\{z\in\mathbb{R}^{K-1}:\ z^{\prime}b_{1}+b_{2}\geq 0\}$.
Since $P(Z\in A_{1})$ and $P(Z\in A_{2})$ are strictly positive, $A_{1}$ and $A_{2}$ are both non-empty. Then there exist $\delta_{1},\delta_{2}>0$ and $z_{1},z_{2}\in\mathcal{Z}$ such that $z_{1}^{\prime}\beta_{1}+\beta_{2}=\delta_{1}$ and $z_{2}^{\prime}b_{1}+b_{2}=-\delta_{2}$. Since $\mathcal{Z}$ is convex and $\mathcal{Z}^{\circ}$ is non-empty, we have that the interior of $\mathcal{Z}$ is dense in $\overline{\mathcal{Z}}$, i.e., $\overline{\mathcal{Z}^{\circ}}=\overline{\mathcal{Z}}$; see Lemma 5.28 of Aliprantis and Border, (2006). By definition, $\mathcal{Z}$ is closed ($\overline{\mathcal{Z}}=\mathcal{Z}$), which means that $\overline{\mathcal{Z}^{\circ}}=\mathcal{Z}$. Thus, we can approximate $z_{1}$ and $z_{2}$ by points in $\mathcal{Z}^{\circ}$. Hence, we can change $\delta_{1},\delta_{2}>0$ slightly such that $z_{1}^{\prime}\beta_{1}+\beta_{2}=\delta_{1}$ and $z_{2}^{\prime}b_{1}+b_{2}=-\delta_{2}$ with $z_{1},z_{2}\in\mathcal{Z}^{\circ}$.
By $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$, we have $z_{1}^{\prime}b_{1}+b_{2}>0$; since $z_{1}\in\overline{A}_{1}\bigcap\mathcal{Z}^{\circ}$, $z_{1}^{\prime}b_{1}+b_{2}\leq 0$ would imply $z_{1}\in\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$ and contradict $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$. Similarly, $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$ implies $z_{2}^{\prime}\beta_{1}+\beta_{2}<0$; since $z_{2}\in\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$, $z_{2}^{\prime}\beta_{1}+\beta_{2}\geq 0$ would imply $z_{2}\in\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$.
Now define $z_{*}(\lambda)=\lambda z_{1}+(1-\lambda)z_{2}$ for $\lambda\in(0,1)$. Since $\mathcal{Z}$ is convex, we have that $\mathcal{Z}^{\circ}$ is also convex by Lemma 5.27 of Aliprantis and Border, (2006). Thus, $z_{*}(\lambda)\in\mathcal{Z}^{\circ}$ for any $\lambda\in(0,1)$. By $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$ and $\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$, we have that $z_{*}(\lambda)\notin A_{1}\bigcap A_{2}$ and $z_{*}(\lambda)\notin B_{1}\bigcap B_{2}$ for any $\lambda\in(0,1)$. This means that $z_{*}(\lambda)^{\prime}\beta_{1}+\beta_{2}$ and $z_{*}(\lambda)^{\prime}b_{1}+b_{2}$ have the same sign. In other words, we have that $f(\lambda)\geq 0$ for any $\lambda\in(0,1)$, where
$f(\lambda)=\left[z_{*}(\lambda)^{\prime}\beta_{1}+\beta_{2}\right]\cdot\left[z_{*}(\lambda)^{\prime}b_{1}+b_{2}\right].$
By straight-forward algebra using $z_{1}^{\prime}\beta_{1}+\beta_{2}=\delta_{1}$ and $z_{2}^{\prime}b_{1}+b_{2}=-\delta_{2}$, we have
$z_{*}(\lambda)^{\prime}\beta_{1}+\beta_{2}=\lambda\delta_{1}+(1-\lambda)\left(z_{2}^{\prime}\beta_{1}+\beta_{2}\right)=\left[\delta_{1}-\left(z_{2}^{\prime}\beta_{1}+\beta_{2}\right)\right]\lambda+\left(z_{2}^{\prime}\beta_{1}+\beta_{2}\right)$ (5)
and
$z_{*}(\lambda)^{\prime}b_{1}+b_{2}=\lambda\left(z_{1}^{\prime}b_{1}+b_{2}\right)-(1-\lambda)\delta_{2}=\left[\left(z_{1}^{\prime}b_{1}+b_{2}\right)+\delta_{2}\right]\lambda-\delta_{2}.$ (6)
Hence, $f(\lambda)$ is a quadratic function of $\lambda$. By $\delta_{1},\delta_{2}>0$, $z_{1}^{\prime}b_{1}+b_{2}>0$ and $z_{2}^{\prime}\beta_{1}+\beta_{2}<0$, the coefficient of the $\lambda^{2}$ term in $f(\lambda)$ satisfies
$\left[\delta_{1}-\left(z_{2}^{\prime}\beta_{1}+\beta_{2}\right)\right]\cdot\left[\left(z_{1}^{\prime}b_{1}+b_{2}\right)+\delta_{2}\right]>\delta_{1}\delta_{2}>0.$
Therefore, $f(\cdot)$ is strictly convex. By (5) and (6), we notice that $f(\lambda_{1})=f(\lambda_{2})=0$, where
$\lambda_{1}=-\frac{z_{2}^{\prime}\beta_{1}+\beta_{2}}{\delta_{1}-\left(z_{2}^{\prime}\beta_{1}+\beta_{2}\right)}\qquad{\rm and}\qquad\lambda_{2}=\frac{\delta_{2}}{\left(z_{1}^{\prime}b_{1}+b_{2}\right)+\delta_{2}}.$
Since $\delta_{1},\delta_{2}>0$, $z_{1}^{\prime}b_{1}+b_{2}>0$ and $z_{2}^{\prime}\beta_{1}+\beta_{2}<0$, we have $\lambda_{1}\in(0,1)$ and $\lambda_{2}\in(0,1)$. We notice that we must have $\lambda_{1}=\lambda_{2}$. Otherwise, there would exist $\lambda$ that is strictly between $\lambda_{1}$ and $\lambda_{2}$, i.e., $\min\{\lambda_{1},\lambda_{2}\}<\lambda<\max\{\lambda_{1},\lambda_{2}\}$. Since $f(\cdot)$ is strictly convex and $f(\lambda_{1})=f(\lambda_{2})=0$, we have that $f(\lambda)<0$ for any $\lambda$ that is strictly between $\lambda_{1}$ and $\lambda_{2}$, contradicting $f(\lambda)\geq 0$ for all $\lambda\in(0,1)$. Hence, we must have that $\lambda_{1}=\lambda_{2}$.
Let $\lambda_{0}$ denote this common value of $\lambda_{1}=\lambda_{2}$. Then $z_{*}(\lambda_{0})^{\prime}\beta_{1}+\beta_{2}=0$ and $z_{*}(\lambda_{0})^{\prime}b_{1}+b_{2}=0$ from (5) and (6). Therefore, we have found $z_{0}:=z_{*}(\lambda_{0})\in\mathcal{Z}^{\circ}$ such that $z_{0}^{\prime}\beta_{1}+\beta_{2}=0$ and $z_{0}^{\prime}b_{1}+b_{2}=0$. In other words, $z_{0}\in\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$ and $z_{0}\in\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}$. This contradicts the assumption that $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$ and $\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}=\emptyset$. ∎
Proof of Theorem 1.
By Lemma 1, Assumption 2 (i.e., $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)>0$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)>0$) implies that $P(Z^{\prime}\beta_{1}+\beta_{2}>0)>0$ and $P(Z^{\prime}\beta_{1}+\beta_{2}<0)>0$.
The rest of the proof proceeds in two steps, in which we prove the “if” and the “only if” parts, respectively.
Step 1: show that if $R(b)=0$, then $b=\mu\beta$ for some $\mu>0$.
We proceed by contradiction. Suppose that $R(b)=0$ and $b\neq\mu\beta$ for any $\mu>0$. Since $R(b)=0$, we have that ${\rm sgn}(E(Y_{1}-Y_{0}\mid X))={\rm sgn}(W^{\prime}b)$ with probability one. By Lemma 1, ${\rm sgn}(E(Y_{1}-Y_{0}\mid X))={\rm sgn}(W^{\prime}\beta)$ with probability one. Hence,
$P(W^{\prime}b<0)=P(E(Y_{1}-Y_{0}\mid X)<0)=P(W^{\prime}\beta<0)$
and
$P(W^{\prime}b>0)=P(E(Y_{1}-Y_{0}\mid X)>0)=P(W^{\prime}\beta>0).$
Recall that we have already proved $P(Z^{\prime}\beta_{1}+\beta_{2}>0)>0$ and $P(Z^{\prime}\beta_{1}+\beta_{2}<0)>0$, i.e., $P(Z\in A_{1})>0$ and $P(Z\in B_{1})>0$. Hence, the above two displays imply that both $P(W^{\prime}b<0)=P(W^{\prime}\beta<0)=P(Z\in B_{1})$ and $P(W^{\prime}b>0)=P(W^{\prime}\beta>0)=P(Z\in A_{1})$ are strictly positive, i.e., $P(Z\in A_{2})>0$ and $P(Z\in B_{2})>0$.
Since both $P(Z\in A_{1})$ and $P(Z\in A_{2})$ are strictly positive, Lemma 6 implies that at least one of $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$ and $\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}$ is non-empty. We discuss theses two cases separately.
Suppose that $\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}$ is non-empty. By the openness of $A_{1}$ and $A_{2}$, $A_{1}\bigcap A_{2}$ is open. We have seen that $P(Z\in A_{1})$, $P(Z\in A_{2})$, $P(Z\in B_{1})$ and $P(Z\in B_{2})$ are all strictly positive. By Lemma 5, we also have that $\overline{A_{1}\bigcap A_{2}}=\overline{A}_{1}\bigcap\overline{A}_{2}$. Thus, $\overline{A_{1}\bigcap A_{2}}\bigcap\mathcal{Z}^{\circ}=\overline{A}_{1}\bigcap\overline{A}_{2}\bigcap\mathcal{Z}^{\circ}\neq\emptyset$. Hence, we apply Lemma 4 with $B=A_{1}\bigcap A_{2}$ and obtain $P(Z\in A_{1}\bigcap A_{2})>0$. By an analogous argument, we have $P(Z\in B_{1}\bigcap B_{2})>0$ if $\overline{B}_{1}\bigcap\overline{B}_{2}\bigcap\mathcal{Z}^{\circ}$ is non-empty.
Therefore, we have proved that at least one of $P(Z\in A_{1}\bigcap A_{2})$ and $P(Z\in B_{1}\bigcap B_{2})$ is strictly positive. Since $R(b)\geq P(A_{1}\bigcap A_{2})+P(B_{1}\bigcap B_{2})$, this means that $R(b)>0$. This contradicts $R(b)=0$. Therefore, we must have that $b=\mu\beta$ for some $\mu>0$.
Step 2: show that if $b=\mu\beta$ for some $\mu>0$, then $R(b)=0$.
Since $b=\mu\beta$ with $\mu>0$, ${\rm sgn}(W^{\prime}b)={\rm sgn}(W^{\prime}\beta)$ almost surely. By Lemma 1, ${\rm sgn}(E(Y_{1}-Y_{0}\mid X))={\rm sgn}(W^{\prime}\beta)$ almost surely. Thus, $R(b)=0$. ∎
A.2 Proof of Corollary 1
Proof of Corollary 1.
Let $d=\dim(X_{t})$. We view this case as a special case of Theorem 1 once we treat $\tilde{X}_{t}=(X_{t}^{\prime},\mathbf{1}\{t=0\})^{\prime}$ as the covariate, where the vector of coefficients is $\tilde{\beta}=(\beta^{\prime},0)^{\prime}$. Since the support of $X_{1}-X_{0}$ contains an open neighborhood of zero, we can restrict the observations to those such that the support of $X_{1}-X_{0}$ is an open neighborhood of zero. Moreover, the sign saturation condition holds because the sign of $E(Y_{1}-Y_{0}\mid X)$ is the same as the sign of $(X_{1}-X_{0})^{\prime}\beta$. By Theorem 1, the vector $(\beta^{\prime},0)^{\prime}$ is identified up to scale. Therefore, $\beta$ is identified up to scale. ∎
A.3 Proof of Theorem 2
Proof of Theorem 2.
We partition $\beta=(\beta_{(1)}^{\prime},\beta_{(2)}^{\prime})^{\prime}$ with $\beta_{(1)}\in\mathbb{R}^{K_{1}}$ and $\beta_{(2)}\in\mathbb{R}^{K_{2}}$. Then $W^{\prime}\beta=D^{\prime}\beta_{(1)}+Z^{\prime}\beta_{(2)}$ and conditional on $D=d_{j}$, we have $W^{\prime}\beta=d_{j}^{\prime}\beta_{(1)}+Z^{\prime}\beta_{(2)}$. Thus, we are back to the situation in which $W=(1,Z^{\prime})^{\prime}$ and the parameter is $(d_{j}^{\prime}\beta_{(1)},\beta_{(2)}^{\prime})^{\prime}$. By the argument of Theorem 1, we can identify $(d_{j}^{\prime}\beta_{(1)},\beta_{(2)}^{\prime})^{\prime}$ up to scale. Clearly, $\beta_{(2)}\neq 0$; otherwise the sign saturation condition on $D=d_{j}$ would fail. By normalizing $\|\beta_{(2)}\|_{2}=1$, we identify $d_{j}^{\prime}\beta_{(1)}$ for each $j\in\{1,...,K_{1}\}$. Since $(d_{1}^{\prime}\beta_{(1)},...,d_{K_{1}}^{\prime}\beta_{(1)})^{\prime}=(d_{1},...,d_{K_{1}})^{\prime}\beta_{(1)}$ and $(d_{1},...,d_{K_{1}})^{\prime}$ has rank $K_{1}$, we identify $\beta_{(1)}$. ∎
A.4 Proof of Theorem 3
In the rest of the appendix, we will use the $p(\cdot,\cdot)$ and $\mathcal{A}(\cdot)$ notations introduced in Definition 2. Recall that for any $t\in\mathbb{R}$, $\mathcal{A}(t)={\rm conv}\{p(t,\alpha):\alpha\in\mathbb{R}\}$ and
$p(t,\alpha)=\begin{pmatrix}F(\alpha)\\ F(t+\alpha)\\ F(\alpha)\cdot F(t+\alpha)\end{pmatrix}.$
We will also repeatedly use the notation $G(a)=\ln\frac{F(a)}{1-F(a)}$ and $\dot{G}(t)=dG(t)/dt$. Throughout the rest of the appendix, we assume that $F(\cdot)$ is strictly increasing on $\mathbb{R}$, which means that $G(\cdot)$ is also strictly increasing on $\mathbb{R}$.
Lemma 7.
Let $s,t\in\mathbb{R}$. Suppose that $\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)>\inf_{\alpha\in\mathbb{R}}v^{\prime}p(t,\alpha)$ for any $v\neq(0,0,0)^{\prime}$. Then $\mathcal{A}(s)\bigcap\mathcal{A}(t)\neq\emptyset$.
Proof.
We proceed by contradiction. Suppose that $\mathcal{A}(s)\bigcap\mathcal{A}(t)=\emptyset$. Notice that both $\mathcal{A}(s)$ and $\mathcal{A}(t)$ are convex and non-empty. By the separating hyperplane theorem (e.g., Theorem 7.30 of Aliprantis and Border, (2006)), there exists $v_{0}\neq(0,0,0)^{\prime}$ and $r\in\mathbb{R}$ such that
$v_{0}^{\prime}a\leq r\qquad\forall a\in\mathcal{A}(s)$
and
$v_{0}^{\prime}b\geq r\qquad\forall b\in\mathcal{A}(t).$
In other words,
[…]
and
[…]
By the assumption and $v_{0}\neq 0$, we have that $\sup_{\alpha}v_{0}^{\prime}p(s,\alpha)>\inf_{\alpha}v_{0}^{\prime}p(t,\alpha)$. Thus, there exists $\varepsilon>0$ such that
$\sup_{\alpha}v_{0}^{\prime}p(s,\alpha)>\varepsilon+\inf_{\alpha}v_{0}^{\prime}p(t,\alpha).$ (7)
[…]
$v_{0}^{\prime}p(t,\xi_{*})\geq v_{0}^{\prime}p(s,\alpha_{*}).$ (8)
By (7) and the definitions of $\alpha_{*}$ and $\xi_{*}$, we have
$v_{0}^{\prime}p(s,\alpha_{*})+\varepsilon/4\geq\sup_{\alpha}v_{0}^{\prime}p(s,\alpha)>\varepsilon+\inf_{\alpha}v_{0}^{\prime}p(t,\alpha)\geq\varepsilon+v_{0}^{\prime}p(t,\xi_{*})-\varepsilon/4.$
This means that $v_{0}^{\prime}p(s,\alpha_{*})\geq\varepsilon/2+v_{0}^{\prime}p(t,\xi_{*})$. Since $\varepsilon>0$, this contradicts (8). ∎
Lemma 8.
Suppose that $s>t$. Then for any $v\in\mathbb{R}^{3}$, $\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)\geq\inf_{\alpha\in\mathbb{R}}v^{\prime}p(t,\alpha)$.
Proof.
Suppose not. Then there exist $v\in\mathbb{R}^{3}$ and $\varepsilon>0$ such that
$\varepsilon+\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)<\inf_{\xi\in\mathbb{R}}v^{\prime}p(t,\xi).$
Let $\delta=s-t$. By assumption, $\delta>0$. The above display implies
$v^{\prime}[p(t+\delta,\alpha)-p(t,\xi)]\leq-\varepsilon\qquad\forall(\alpha,\xi)\in\mathbb{R}^{2}.$
Let $v=(v_{1},v_{2},v_{3})^{\prime}$. Then for any $\alpha,\xi\in\mathbb{R}$,
$v_{1}[F(\alpha)-F(\xi)]+v_{2}[F(t+\delta+\alpha)-F(t+\xi)]+v_{3}[F(\alpha)F(t+\delta+\alpha)-F(\xi)F(t+\xi)]\leq-\varepsilon.$ (9)
Taking $\xi=\alpha$, we obtain that for any $\alpha\in\mathbb{R}$,
$v_{2}[F(t+\delta+\alpha)-F(t+\alpha)]+v_{3}F(\alpha)[F(t+\delta+\alpha)-F(t+\alpha)]\leq-\varepsilon.$
This means that for any $\alpha\in\mathbb{R}$,
$v_{2}\leq-v_{3}F(\alpha)-\frac{\varepsilon}{F(t+\delta+\alpha)-F(t+\alpha)}.$
Notice that
$\inf_{\alpha\in\mathbb{R}}\left[-v_{3}F(\alpha)-\frac{\varepsilon}{F(t+\delta+\alpha)-F(t+\alpha)}\right]\leq\inf_{\alpha\in\mathbb{R}}\left[|v_{3}|-\frac{\varepsilon}{F(t+\delta+\alpha)-F(t+\alpha)}\right]\\ \leq\lim_{\alpha\rightarrow-\infty}\left[|v_{3}|-\frac{\varepsilon}{F(t+\delta+\alpha)-F(t+\alpha)}\right]=-\infty.$
Thus, $v_{2}\leq-\infty$. This contradicts $v\in\mathbb{R}^{3}$. ∎
Lemma 9.
Assume that $s>t>0$. If $\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)=\inf_{\alpha\in\mathbb{R}}v^{\prime}p(t,\alpha)$ for some $v\neq(0,0,0)^{\prime}$, then
$\inf_{\alpha\in\mathbb{R}}[G(s+\alpha)-G(\alpha)]\geq\sup_{\xi\in\mathbb{R}}[G(t+\xi)-G(\xi)].$
Proof.
Let $v=(v_{1},v_{2},v_{3})^{\prime}$. Without loss of generality, we can normalize and assume $v_{3}\in\{0,1,-1\}$. Notice that $\lim_{\alpha\rightarrow\infty}v^{\prime}p(t,\alpha)=0$ and $\lim_{\alpha\rightarrow\infty}v^{\prime}p(s,\alpha)=0$. Thus, $\inf_{\alpha\in\mathbb{R}}v^{\prime}p(t,\alpha)\leq 0\leq\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)$. By $\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)=\inf_{\alpha\in\mathbb{R}}v^{\prime}p(t,\alpha)$, it follows that
$\sup_{\alpha\in\mathbb{R}}v^{\prime}p(s,\alpha)=\inf_{\alpha\in\mathbb{R}}v^{\prime}p(t,\alpha)=0.$
Let $\delta=s-t$. By assumption, $\delta>0$ and $t>0$. Then for any $\alpha,\xi\in\mathbb{R}$, $v^{\prime}p(s,\alpha)\leq 0\leq v^{\prime}p(t,\xi)$. In other words, for any $\alpha,\xi\in\mathbb{R}$,
$v_{1}F(\alpha)+v_{2}F(t+\delta+\alpha)+v_{3}F(\alpha)F(t+\delta+\alpha)\leq 0\\ \leq v_{1}F(\xi)+v_{2}F(t+\xi)+v_{3}F(\xi)F(t+\xi).$ (10)
We now proceed in three steps.
Step 1: rule out $v_{3}=0$.
Suppose that $v_{3}=0$. By (10), we have that for any $\alpha,\xi\in\mathbb{R}$,
$v_{1}[F(\alpha)-F(\xi)]+v_{2}[F(t+\delta+\alpha)-F(t+\xi)]\leq 0.$ (11)
Take $\alpha=\xi-\delta$ and obtain $v_{1}[F(\xi-\delta)-F(\xi)]\leq 0$ for any $\xi\in\mathbb{R}$. By $\delta>0$, this means that $v_{1}\geq 0$.
It is easy to verify $v_{1}+v_{2}=0$: taking $\alpha=-\infty$ in (11), we have that $-v_{1}F(\xi)-v_{2}F(t+\xi)\leq 0$ for any $\xi\in\mathbb{R}$, which means that $v_{1}+v_{2}\geq 0$ (take $\xi\rightarrow\infty$); taking $\alpha=\infty$ in (11), we have that $v_{1}(1-F(\xi))+v_{2}(1-F(t+\xi))\leq 0$ for any $\xi\in\mathbb{R}$, which means that $v_{1}+v_{2}\leq 0$ (take $\xi\rightarrow-\infty$).
Hence, $v=(v_{1},-v_{1},0)$. By $v\neq(0,0,0)^{\prime}$ and $v_{1}\geq 0$, we have $v_{1}>0$. By (11), we have that for any $\alpha,\xi\in\mathbb{R}$,
$v_{1}[F(\alpha)-F(\xi)]-v_{1}[F(t+\delta+\alpha)-F(t+\xi)]\leq 0.$
Since $v_{1}>0$, it follows that for any $\alpha,\xi\in\mathbb{R}$,
$F(\alpha)-F(\xi)\leq F(t+\delta+\alpha)-F(t+\xi).$
Taking $\alpha=\infty$, we have $F(\xi)\geq F(t+\xi)$, which is impossible because $t>0$. Thus, $v_{3}\neq 0$.
Step 2: rule out $v_{3}=1$.
Suppose that $v_{3}=1$. Then (10) implies that for any $\alpha,\xi\in\mathbb{R}$,
$v_{1}[F(\alpha)-F(\xi)]+v_{2}[F(t+\delta+\alpha)-F(t+\xi)]+F(\alpha)F(t+\delta+\alpha)-F(\xi)F(t+\xi)\leq 0.$ (12)
Taking $\alpha\rightarrow\infty$ and $\xi\rightarrow-\infty$, we have $v_{1}+v_{2}+1\leq 0$. Taking $\alpha\rightarrow-\infty$ and $\xi\rightarrow\infty$, we have $-v_{1}-v_{2}-1\leq 0$, which means $v_{1}+v_{2}+1\geq 0$. Hence, $v_{1}+v_{2}+1=0$.
Take $\alpha=\xi-\delta$ in (12) and obtain that for any $\xi\in\mathbb{R}$,
$v_{1}[F(\xi-\delta)-F(\xi)]+[F(\xi-\delta)-F(\xi)]F(t+\xi)\leq 0.$
Since $\delta>0$, we have $v_{1}+F(t+\xi)\geq 0$ for any $\xi\in\mathbb{R}$. Taking $\xi\rightarrow-\infty$, we have $v_{1}\geq 0$.
By $v_{3}=1$ and $v_{1}+v_{2}+1=0$, the right-hand side of (10) implies that for any $\xi\in\mathbb{R}$,
$v_{1}F(\xi)+(-1-v_{1})F(t+\xi)+F(\xi)F(t+\xi)\geq 0.$
Taking $\xi=0$, we have $[F(0)-1]F(t)\geq v_{1}[F(t)-F(0)]$. Since $t>0$, we have
$v_{1}\leq\frac{[F(0)-1]F(t)}{F(t)-F(0)}<0.$
This contradicts $v_{1}\geq 0$. Thus, $v_{3}\neq 1$.
Step 3: show the final result.
From the previous two steps, we ruled out $v_{3}\in\{0,1\}$. Hence, $v_{3}=-1$. Therefore, (10) implies that for any $\alpha,\xi\in\mathbb{R}$,
$[v_{1}-F(t+\delta+\alpha)][F(\alpha)-F(\xi)]+[v_{2}-F(\xi)][F(t+\delta+\alpha)-F(t+\xi)]\leq 0$ (13)
Taking $\alpha=\xi-\delta$ gives $v_{1}-F(t+\xi)\geq 0$ (due to $\delta>0$) for any $\xi\in\mathbb{R}$, which means that $v_{1}\geq 1$.
By the right-hand side of (10) and $v_{3}=-1$, $F(\xi)v_{1}+v_{2}F(t+\xi)-F(\xi)F(t+\xi)\geq 0$ for any $\xi\in\mathbb{R}$. Taking $\xi=\infty$ gives $v_{1}+v_{2}\geq 1$. On the other hand, the left-hand side of (10) and $v_{3}=-1$ yield $v_{1}F(\alpha)+v_{2}F(t+\delta+\alpha)-F(\alpha)F(t+\delta+\alpha)\leq 0$ for any $\alpha\in\mathbb{R}$, which (by taking $\alpha\rightarrow\infty$) yields $v_{1}+v_{2}\leq 1$. Hence, $v_{1}+v_{2}=1$, which means that $v_{2}=1-v_{1}$ and $v=(v_{1},1-v_{1},-1)^{\prime}$.
Using the left-hand side of (10) and $v=(v_{1},1-v_{1},-1)^{\prime}$, we have that for any $\alpha\in\mathbb{R}$,
$v_{1}F(\alpha)+(1-v_{1})F(t+\delta+\alpha)-F(\alpha)F(t+\delta+\alpha)\leq 0.$
Since $F(t+\delta+\alpha)>F(\alpha)$ (due to $t+\delta>0$), we have that for any $\alpha\in\mathbb{R}$,
$v_{1}\geq\frac{F(t+\delta+\alpha)[1-F(\alpha)]}{F(t+\delta+\alpha)-F(\alpha)}.$
Hence,
$v_{1}\geq\sup_{\alpha\in\mathbb{R}}\frac{F(t+\delta+\alpha)[1-F(\alpha)]}{F(t+\delta+\alpha)-F(\alpha)}.$ (14)
Similarly, the right-hand side of (10) and $v=(v_{1},1-v_{1},-1)^{\prime}$, we have that for any $\xi\in\mathbb{R}$, $v_{1}F(\xi)+(1-v_{1})F(t+\xi)-F(\xi)F(t+\xi)\geq 0$. Hence,
$v_{1}\leq\inf_{\xi\in\mathbb{R}}\frac{F(t+\xi)[1-F(\xi)]}{F(t+\xi)-F(\xi)}.$ (15)
Combining (14) and (15), we have
$\sup_{\alpha\in\mathbb{R}}\frac{F(t+\delta+\alpha)[1-F(\alpha)]}{F(t+\delta+\alpha)-F(\alpha)}\leq\inf_{\xi\in\mathbb{R}}\frac{F(t+\xi)[1-F(\xi)]}{F(t+\xi)-F(\xi)}.$ (16)
Recall $G(a)=\ln\frac{F(a)}{1-F(a)}$. Then with straight-forward algebra, (16) becomes $\inf_{\alpha\in\mathbb{R}}\left[G(t+\delta+\alpha)-G(\alpha)\right]\geq\sup_{\xi\in\mathbb{R}}\left[G(t+\xi)-G(\xi)\right]$. ∎
Lemma 10.
Let $T\subset(0,\infty)$ be a compact set. Assume that $G(\cdot)$ is continuously differentiable. Suppose that for any $\delta>0$, there exists $t\in T$ such that
$\inf_{\alpha\in\mathbb{R}}[G(t+\delta+\alpha)-G(\alpha)]\geq\sup_{\xi\in\mathbb{R}}[G(t+\xi)-G(\xi)].$
Then $\dot{G}(\cdot)$ is a periodic function with a period in $T$. In other words, there exists $t\in T$ such that $\dot{G}(t+\xi)=\dot{G}(\xi)$ for any $\xi\in\mathbb{R}$.
Proof.
For any $b\in\mathbb{R}$, define $r_{b}(\cdot)$ by $r_{b}(a)=G(b+a)-G(a)$. We also define $c(b)=\sup_{\xi\in\mathbb{R}}r_{b}(\xi)$. Let $K>0$ be arbitrary. Then by assumption, for any $\delta>0$, there exists $t\in T$ (depending only on $\delta$) such that $\inf_{\alpha\in\mathbb{R}}[r_{t}(\delta+\alpha)+r_{\delta}(\alpha)]\geq\sup_{\xi\in\mathbb{R}}r_{t}(\xi)=c(t)$, which means that $\inf_{\alpha+\delta\in[-K,K]}[r_{t}(\delta+\alpha)+r_{\delta}(\alpha)]\geq c(t)$ and thus
$c(t)\leq\inf_{|\delta+\alpha|\leq K}r_{t}(\delta+\alpha)+\sup_{|\delta+\alpha|\leq K}r_{\delta}(\alpha)=\inf_{|\xi|\leq K}r_{t}(\xi)+\sup_{|\delta+\alpha|\leq K}r_{\delta}(\alpha).$
Notice that $\sup_{|\delta+\alpha|\leq K}r_{\delta}(\alpha)=\sup_{|\delta+\alpha|\leq K}[G(\delta+\alpha)-G(\alpha)]=\sup_{|\xi|\leq K}[G(\xi)-G(\xi-\delta)]$. We have that for any $\delta,K>0$, there exists $t\in T$ depending only on $\delta$ such that
$\sup_{|\xi|\leq K}[G(\xi)-G(\xi-\delta)]\geq c(t)-\inf_{|\xi|\leq K}r_{t}(\xi)=\sup_{\xi\in\mathbb{R}}r_{t}(\xi)-\inf_{|\xi|\leq K}r_{t}(\xi)\geq 0.$
Now we choose $\delta=1/n$. Then for any $n>1$, there exists $t_{n}$ (not depending on $K$) such that
$\sup_{|\xi|\leq K}[G(\xi)-G(\xi-1/n)]\geq c(t_{n})-\inf_{|\xi|\leq K}r_{t_{n}}(\xi)\geq 0.$
Let $q_{n}(K):=c(t_{n})-\inf_{|\xi|\leq K}r_{t_{n}}(\xi)$. Since $G(\cdot)$ is continuous, it is uniformly continuous on compact sets (Heine-Cantor theorem). Hence, $\limsup_{n\rightarrow\infty}\sup_{|\xi|\leq K}[G(\xi)-G(\xi-1/n)]=0$. The above display implies $\limsup_{n\rightarrow\infty}q_{n}(K)=0$.
Since $t_{n}$ is in a compact set $T$, there exists a subsequence $t_{n_{j}}$ and $t_{*}\in T$ such that $t_{n_{j}}\rightarrow t_{*}$ and $t_{*}$ does not depend on $K$. We can further extract a subsequence that is either increasing or decreasing. With an abuse of notation, we write $t_{n}$ rather than $t_{n_{j}}$ in the rest of the proof. We show the following claim in two cases ($t_{n}\uparrow t_{*}$ and $t_{n}\downarrow t_{*}$):
$\sup_{|\xi|\leq K}r_{t_{*}}(\xi)=\inf_{|\xi|\leq K}r_{t_{*}}(\xi).$ (17)
Step 1: show (17) assuming $t_{n}\uparrow t_{*}$.
Since $t_{n}\leq t_{*}$, we have $r_{t_{n}}(\xi)\leq r_{t_{*}}(\xi)$ for any $\xi$ and thus
$c(t_{n})=q_{n}+\inf_{|\xi|\leq K}r_{t_{n}}(\xi)\leq q_{n}+\inf_{|\xi|\leq K}r_{t_{*}}(\xi).$
Hence, by $\limsup_{n\rightarrow\infty}q_{n}(K)=0$, we have
$\limsup_{n\rightarrow\infty}c(t_{n})\leq\inf_{|\xi|\leq K}r_{t_{*}}(\xi).$ (18)
Let $R=\max_{t\in T}|t|$. Since $T$ is compact, $R$ is bounded. Notice that $G(\cdot)$ is continuous. Again, by the Heine-Cantor theorem, it is uniformly continuous on the compact set $[-K-R,K+R]$. Notice that $t_{n}+\xi\in[-K-R,K+R]$ for any $\xi\in[-K,K]$ (due to $|t_{n}|\leq R$). Thus, $\sup_{|\xi|\leq K}|r_{t_{n}}(\xi)-r_{t_{*}}(\xi)|=\sup_{|\xi|\leq K}|G(t_{n}+\xi)-G(t_{*}+\xi)|\rightarrow 0.$ It follows that
$\limsup_{n\rightarrow\infty}\sup_{|\xi|\leq K}r_{t_{n}}(\xi)=\sup_{|\xi|\leq K}r_{t_{*}}(\xi).$
Now by (18) and the above display, we have that
$\inf_{|\xi|\leq K}r_{t_{*}}(\xi)\geq\limsup_{n\rightarrow\infty}c(t_{n})=\limsup_{n\rightarrow\infty}\sup_{\xi\in\mathbb{R}}r_{t_{n}}(\xi)\geq\limsup_{n\rightarrow\infty}\sup_{|\xi|\leq K}r_{t_{n}}(\xi)=\sup_{|\xi|\leq K}r_{t_{*}}(\xi).$
On the other hand, we have $\inf_{|\xi|\leq K}r_{t_{*}}(\xi)\leq\sup_{|\xi|\leq K}r_{t_{*}}(\xi)$. This proves (17).
Step 2: show (17) assuming $t_{n}\downarrow t_{*}$.
Since the argument is similar to Step 1, we only provide an outline here. By $t_{n}\geq t_{*}$, we have
$q_{n}+\inf_{|\xi|\leq K}r_{t_{n}}(\xi)=c(t_{n})\geq c(t_{*})=\sup_{\xi\in\mathbb{R}}r_{t_{*}}(\xi)\geq\sup_{|\xi|\leq K}r_{t_{*}}(\xi).$
Notice that $r_{t_{*}}(\xi)=G(t_{*}+\xi)-G(\xi)$ and $r_{t_{n}}(\xi)=G(t_{n}+\xi)-G(\xi)$. Again, by the Heine-Cantor theorem and the continuity of $G(\cdot)$, we have $\sup_{|\xi|\leq K}|r_{t_{n}}(\xi)-r_{t_{*}}(\xi)|=\sup_{|\xi|\leq K}|G(t_{n}+\xi)-G(t_{*}+\xi)|\rightarrow 0$ due to $t_{n}\rightarrow t_{*}$. Thus,
$\liminf_{n\rightarrow\infty}\inf_{|\xi|\leq K}r_{t_{n}}(\xi)=\inf_{|\xi|\leq K}r_{t_{*}}(\xi).$
The above two displays and $q_{n}\rightarrow 0$ imply $\inf_{|\xi|\leq K}r_{t_{*}}(\xi)\geq\sup_{|\xi|\leq K}r_{t_{*}}(\xi)$. This gives (17).
Step 3: show the final result.
Notice that we have verified (17) for an arbitrary $K>0$. Thus, $r_{t_{*}}(\cdot)$ is a constant function on $[-K,K]$ for any $K>0$. Since $t_{*}$ does not depend on $K$, $r_{t_{*}}(\cdot)$ is a constant function on $\mathbb{R}$. This means that the derivative of $r_{t_{*}}(\cdot)$ is zero on $\mathbb{R}$. Recalling that $r_{t_{*}}(\xi)=G(t_{*}+\xi)-G(\xi)$, we have $\dot{G}(t_{*}+\xi)=\dot{G}(\xi)$ for any $\xi\in\mathbb{R}$. ∎
Proof of Theorem 3.
Fix an arbitrary $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}\in\mathcal{B}_{+}$. Then $w^{\prime}\beta=z^{\prime}\beta_{1}+\beta_{2}>0$ for any $z\in\mathcal{Z}$. Let $\delta>0$ be a number to be determined. Define $b=(\beta_{1}^{\prime},\beta_{2}+\delta)^{\prime}$. Clearly, $w^{\prime}b=z^{\prime}\beta_{1}+\beta_{2}+\delta>0$ for any $z\in\mathcal{Z}$. Thus, $b\in\mathcal{B}_{+}$. Define $T$ to be the closure of $\{z^{\prime}\beta_{1}+\beta_{2}:\ z\in\mathcal{Z}\}$. Since $\mathcal{Z}$ is bounded, $T$ is also bound. By the closedness of $T$, $T$ is also compact. We also notice that $0\notin T$ and thus $T\subset(0,\infty)$. (To see this, notice that $\inf_{z\in\mathcal{Z}}(z^{\prime},1)^{\prime}\beta$ is achieved by some $z_{*}\in\mathcal{Z}$ (because $\mathcal{Z}$ is closed and bounded). On the other hand, since for any $z\in\mathcal{Z}$, $(z^{\prime},1)^{\prime}\beta>0$, we have that $(z_{*}^{\prime},1)^{\prime}\beta>0$.)
We now show that there exists $\delta>0$ such that for any $z\in\mathcal{Z}$, $\mathcal{A}(w^{\prime}\beta)\bigcap\mathcal{A}(w^{\prime}b)\neq\emptyset$, where $w=(z^{\prime},1)^{\prime}$. We proceed by contradiction. Suppose that for any $\delta>0$, there exists $z_{\delta}\in\mathcal{Z}$ such that $\mathcal{A}(w_{\delta}^{\prime}b)\bigcap\mathcal{A}(w_{\delta}^{\prime}\beta)=\emptyset$, where $w_{\delta}=(z_{\delta}^{\prime},1)^{\prime}$. By Lemma 7, there exists $v_{\delta}\neq 0$ such that $\sup_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}b,\alpha)\leq\inf_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}\beta,\alpha)$. By Lemma 8 and $w_{\delta}^{\prime}b-w_{\delta}^{\prime}\beta=\delta>0$, we have that $\sup_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}b,\alpha)\geq\inf_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}\beta,\alpha)$. Hence,
$\sup_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}\beta+\delta,\alpha)=\sup_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}b,\alpha)=\inf_{\alpha\in\mathbb{R}}v_{\delta}^{\prime}p(w_{\delta}^{\prime}\beta,\alpha).$
By Lemma 9 (with $t=w_{\delta}^{\prime}\beta$),
$\inf_{\alpha\in\mathbb{R}}[G(w_{\delta}^{\prime}\beta+\delta+\alpha)-G(\alpha)]\geq\sup_{\xi\in\mathbb{R}}[G(w_{\delta}^{\prime}\beta+\xi)-G(\xi)].$
Notice that $w_{\delta}^{\prime}\beta\in T$. Therefore, we have shown that for any $\delta>0$, there exists $t_{\delta}\in T$ such that
$\inf_{\alpha\in\mathbb{R}}[G(t_{\delta}+\delta+\alpha)-G(\alpha)]\geq\sup_{\xi\in\mathbb{R}}[G(t_{\delta}+\xi)-G(\xi)].$
By Lemma 10, $\dot{G}$ is a periodic function. However, this contradicts the assumption that $\dot{G}$ is not a periodic function. ∎
A.5 Proof of Theorem 4
Lemma 11.
Suppose that $h(\cdot)$ is a continuous and non-constant function on $\mathbb{R}$. If $h(\cdot)$ is a periodic function, then $h(\cdot)$ has a minimal positive period, i.e., the set $\{a>0:\ h(a+x)=h(x)\ \forall x\in\mathbb{R}\}$ has a smallest element.
Proof.
Let $K=\{a>0:\ h(a+x)=h(x)\ \forall x\in\mathbb{R}\}$.
We first show that $\inf K>0$. Suppose that this is not true. Then $\inf K=0$. Thus, there exists $a_{n}\in K$ such that $a_{n}\downarrow 0$. Let $x,y\in\mathbb{R}$ such that $x<y$. For any $n$, define $J_{n}$ to be the integer part of $(y-x)/a_{n}$, i.e., $J_{n}$ is the integer satisfying $J_{n}a_{n}\leq y-x<(J_{n}+1)a_{n}$. Since $a_{n}$ is a period for $h(\cdot)$ and $J_{n}$ is an integer, $J_{n}a_{n}$ is also a period of $h(\cdot)$. Thus, $h(y)=h(y-x+x)=h(y-x-J_{n}a_{n}+x)$. Notice that $0\leq y-x-J_{n}a_{n}<a_{n}\rightarrow 0$. Thus, $(y-x-J_{n}a_{n})+x\rightarrow x$. By the continuity of $h(\cdot)$, we have that $h(y-x-J_{n}a_{n}+x)\rightarrow h(x)$. Hence, we have proved that $h(y)=h(x)$. Since $x,y$ are two arbitrary numbers with $x<y$, this means that $h(\cdot)$ is a constant function on $\mathbb{R}$. This would contradict the assumption that $h(\cdot)$ is non-constant. Hence, $\inf K>0$.
Let $\eta=\inf K$. Since $\eta>0$ and $\eta\leq a$ for any $a\in K$, we only need to show that $\eta\in K$. By definition, there exists $a_{n}\in K$ with $a_{n}\rightarrow\eta$. For any $x$, we apply the continuity of $h(\cdot)$ and obtain that $h(x)=h(x+a_{n})\rightarrow h(x+\eta)$, which means that $h(x)=h(x+\eta)$ and thus $\eta\in K$. Therefore, $\eta$ is the smallest element of $K$. ∎
Lemma 12.
Define the function $G(a)=\ln\frac{F(a)}{1-F(a)}$. Assume that $s>t>0$. If
$\inf_{\alpha\in\mathbb{R}}[G(s+\alpha)-G(\alpha)]>\sup_{\xi\in\mathbb{R}}[G(t+\xi)-G(\xi)]>0,$
then $\mathcal{A}(s)\bigcap\mathcal{A}(t)=\emptyset$.
Proof.
Let $\delta=s-t$. By $s>t$, $\delta>0$. We proceed in two steps.
Step 1: show that there exists $v\in\mathbb{R}^{3}$ such that $v^{\prime}p(s,\alpha)<0<v^{\prime}p(t,\xi)$ for any $\alpha,\xi\in\mathbb{R}$.
Consider $v=(v_{1},1-v_{1},-1)^{\prime}$. We would like to choose $v_{1}$ such that for any $\alpha,\xi\in\mathbb{R}$
$v_{1}F(\alpha)+(1-v_{1})F(s+\alpha)-F(\alpha)F(s+\alpha)<0$ (19)
and
$v_{1}F(\xi)+(1-v_{1})F(t+\xi)-F(\xi)F(t+\xi)>0.$ (20)
Since $s>t>0$, it suffices to choose any $v_{1}$ such that
$\sup_{\alpha\in\mathbb{R}}\frac{F(s+\alpha)[1-F(\alpha)]}{F(s+\alpha)-F(\alpha)}<v_{1}<\inf_{\xi\in\mathbb{R}}\frac{F(t+\xi)[1-F(\xi)]}{F(t+\xi)-F(\xi)}.$
By $F(\cdot)=[\exp(-G(\cdot))+1]^{-1}$, this means that we can choose any $v_{1}$ such that
$\frac{1}{1-\sup_{\alpha\in\mathbb{R}}\exp[G(\alpha)-G(s+\alpha)]}<v_{1}<\frac{1}{1-\inf_{\xi\in\mathbb{R}}\exp[G(\xi)-G(t+\xi)]}.$
Since $\inf_{\alpha\in\mathbb{R}}[G(s+\alpha)-G(\alpha)]>\sup_{\xi\in\mathbb{R}}[G(t+\xi)-G(\xi)]>0$, we have $\sup_{\alpha\in\mathbb{R}}[G(\alpha)-G(s+\alpha)]<\inf_{\xi\in\mathbb{R}}[G(\xi)-G(t+\xi)]<0$ and thus a choice of $v_{1}$ in the above display is clearly possible. Therefore, by $v=(v_{1},1-v_{1},-1)^{\prime}$, we have found $v\in\mathbb{R}^{3}$ such that (19) and (20) hold.
Step 2: show the final result.
We proceed by contradiction. Suppose that $\mathcal{A}(s)\bigcap\mathcal{A}(t)\neq\emptyset$. Then there exists $r\in\mathcal{A}(s)\bigcap\mathcal{A}(t)$. Since $r\in\mathcal{A}(s)\subset\mathbb{R}^{3}$, by Caratheodory’s theorem (e.g., Theorem 5.32 in Aliprantis and Border, (2006)), there exist $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}\in\mathbb{R}$ and $\lambda_{1},...,\lambda_{4}\geq 0$ such that $r=\sum_{j=1}^{4}\lambda_{j}p(s,\alpha_{j})$ and $\sum_{j=1}^{4}\lambda_{j}=1$. Similarly, $r=\sum_{j=1}^{4}\rho_{j}p(t,\xi_{j})$, where $\xi_{1},...,\xi_{4}\in\mathbb{R}$ and $\rho_{1},...,\rho_{4}\geq 0$ satisfy $\sum_{j=1}^{4}\rho_{j}=1$. Therefore, for any $v\in\mathbb{R}^{3}$,
$\max_{1\leq j\leq 4}p(s,\alpha_{j})^{\prime}v\geq\sum_{j=1}^{4}\lambda_{j}p(s,\alpha_{j})^{\prime}v=r^{\prime}v=\sum_{j=1}^{4}\rho_{j}p(t,\xi_{j})^{\prime}v\geq\min_{1\leq j\leq 4}p(t,\xi_{j})^{\prime}v.$
However, in Step 1, we have shown that there exists $v\neq(0,0,0)^{\prime}$ such that $v^{\prime}p(s,\alpha)<0<v^{\prime}p(t,\xi)$ for any $\alpha,\xi\in\mathbb{R}$. This is a contradiction. ∎
Lemma 13.
Assume that $s>t>0$. Suppose that $\dot{G}$ is a periodic function with a positive period $\eta>0$. If either $s/\eta$ or $t/\eta$ is an integer, then there exists $\varepsilon\in(0,(s-t)/4]$ such that
$\inf_{\alpha\in\mathbb{R}}[G(s-\varepsilon+\alpha)-G(\alpha)]>\sup_{\xi\in\mathbb{R}}[G(t+\varepsilon+\xi)-G(\xi)]>0.$
Proof.
Since $\dot{G}(l\cdot\eta+\xi)-\dot{G}(\xi)=0$ for any $\xi\in\mathbb{R}$ and for any $l\in\mathbb{Z}$ ($\mathbb{Z}$ denotes the set of all integers), it follows that $G(l\cdot\eta+\xi)-G(\xi)$ does not depend on $\xi$, which means that $G(l\cdot\eta+\xi)-G(\xi)=G(l\cdot\eta+0)-G(0)$. We notice that
[…]
where $q_{0}=G(\eta)-G(0)$. Thus, we have that for any $\xi\in\mathbb{R}$ and for any $l\in\mathbb{Z}$,
$G(l\cdot\eta+\xi)=G(\xi)+lq_{0}.$ (21)
Notice that we can represent all the real numbers as $(k+r)\eta$ with $k\in\mathbb{Z}$ and $r\in[0,1)$. Then it suffices to show that for some $\varepsilon>0$,
$\inf_{(k_{1},r_{1})\in\mathbb{Z}\times[0,1)}[G(s-\varepsilon+(k_{1}+r_{1})\eta)-G((k_{1}+r_{1})\eta)]\\ >\sup_{(k_{2},r_{2})\in\mathbb{Z}\times[0,1)}[G(t+\varepsilon+(k_{2}+r_{2})\eta)-G((k_{2}+r_{2})\eta)]>0.$
By (21), $G(s-\varepsilon+(k_{1}+r_{1})\eta)-G((k_{1}+r_{1})\eta)=G(s-\varepsilon+r_{1}\eta)-G(r_{1}\eta)$, which does not depend on $k_{1}$; similarly, $G(t+\varepsilon+(k_{2}+r_{2})\eta)-G((k_{2}+r_{2})\eta)=G(t+\varepsilon+r_{2}\eta)-G(r_{2}\eta)$. Thus, it suffices to show that for some $\varepsilon\in(0,(s-t)/4]$,
$\inf_{r_{1}\in[0,1)}[G(s-\varepsilon+r_{1}\eta)-G(r_{1}\eta)]>\sup_{r_{2}\in[0,1)}[G(t+\varepsilon+r_{2}\eta)-G(r_{2}\eta)].$ (22)
We now construct $\varepsilon$ that satisfies (22). We consider two cases: $s/\eta\in\mathbb{Z}$ or $t/\eta\in\mathbb{Z}$.
Case 1: suppose $s/\eta\in\mathbb{Z}$. Let $k_{1}=s/\eta$. Then by (21), we have $G(s+r_{1}\eta)-G(r_{1}\eta)=k_{1}q_{0}$ and $G(t+r_{2}\eta)=G(t-s+r_{2}\eta)+k_{1}q_{0}$. Hence,
$\inf_{r_{1}\in[0,1)}[G(s+r_{1}\eta)-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(t+r_{2}\eta)-G(r_{2}\eta)]$
$=k_{1}q_{0}-\sup_{r_{2}\in[0,1)}[G(t-s+r_{2}\eta)+k_{1}q_{0}-G(r_{2}\eta)]$
$=\inf_{r_{2}\in[0,1)}[G(r_{2}\eta)-G(r_{2}\eta-(s-t))]\geq\inf_{r_{2}\in[0,1]}[G(r_{2}\eta)-G(r_{2}\eta-(s-t))].$
Since $r_{2}\mapsto G(t+r_{2}\eta)-G(t-s+r_{2}\eta)$ is a continuous function and $[0,1]$ is a compact set, the above infimum is achieved at point in $[0,1]$. Since $G$ is strictly increasing and $s-t>0$, we have that
$\inf_{r_{2}\in[0,1]}[G(r_{2}\eta)-G(r_{2}\eta-(s-t))]>0.$
(Otherwise, there would exist $r_{2,*}\in[0,1]$ such that $r_{2,*}\eta=r_{2,*}\eta-(s-t)$.) Let $\Delta_{1}=\inf_{r_{2}\in[0,1]}[G(r_{2}\eta)-G(r_{2}\eta-(s-t))]$. We now choose $\varepsilon\in(0,(s-t)/4]$ such that
$\sup_{r_{1}\in[0,1]}[G(r_{1}\eta)-G(r_{1}\eta-\varepsilon)]\leq\Delta_{1}/4$
and
$\sup_{r_{2}\in[0,1]}[G(r_{2}\eta-(s-t)+\varepsilon)-G(r_{2}\eta-(s-t))]\leq\Delta_{1}/4$
To see that this is possible, notice that $G(\cdot)$ is continuous and thus is uniformly continuous on the compact set $[-1,\eta]\bigcup[-(s-t),\eta+1-(s-t)]$. Thus, there exists a constant $\kappa_{1}>0$ such that $\sup_{r_{1}\in[0,1]}[G(r_{1}\eta)-G(r_{1}\eta-\varepsilon)]\leq\kappa_{1}\varepsilon$ and $\sup_{r_{2}\in[0,1]}[G(r_{2}\eta-(s-t)+\varepsilon)-G(r_{2}\eta-(s-t))]\leq\kappa_{1}\varepsilon$. Hence, we can simply choose $\varepsilon=\min\{(s-t)/4,\Delta_{1}/(4\kappa_{1})\}$. Therefore,
$\inf_{r_{1}\in[0,1)}[G(s-\varepsilon+r_{1}\eta)-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(t+\varepsilon+r_{2}\eta)-G(r_{2}\eta)]$
$=\inf_{r_{1}\in[0,1)}[G(-\varepsilon+r_{1}\eta)+k_{1}q_{0}-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(t-s+\varepsilon+r_{2}\eta)+k_{1}q_{0}-G(r_{2}\eta)]$
$=\inf_{r_{1}\in[0,1)}[G(-\varepsilon+r_{1}\eta)-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(t-s+\varepsilon+r_{2}\eta)-G(r_{2}\eta)]$
$=-\sup_{r_{1}\in[0,1)}[G(r_{1}\eta)-G(r_{1}\eta-\varepsilon)]+\inf_{r_{2}\in[0,1)}[G(r_{2}\eta)-G(r_{2}\eta-(s-t)+\varepsilon)]$
$\geq-\sup_{r_{1}\in[0,1)}[G(r_{1}\eta)-G(r_{1}\eta-\varepsilon)]+\inf_{r_{2}\in[0,1)}[G(r_{2}\eta)-G(r_{2}\eta-(s-t))]$
$\qquad+\inf_{r_{2}\in[0,1)}[G(r_{2}\eta-(s-t))-G(r_{2}\eta-(s-t)+\varepsilon)]$
$=-\sup_{r_{1}\in[0,1)}[G(r_{1}\eta)-G(r_{1}\eta-\varepsilon)]+\inf_{r_{2}\in[0,1)}[G(r_{2}\eta)-G(r_{2}\eta-(s-t))]$
$\qquad-\sup_{r_{2}\in[0,1)}[G(r_{2}\eta-(s-t)+\varepsilon)-G(r_{2}\eta-(s-t))]$
[…]
where (i) follows by the construction of $\varepsilon$ and (ii) follows by the definition of $\Delta_{1}$.
Case 2: suppose $t/\eta\in\mathbb{Z}$. Let $k_{2}=t/\eta$. Then by (21), we have $G(t+r_{2}\eta)-G(r_{2}\eta)=k_{2}q_{0}$ and $G(s+r_{1}\eta)=G(s-t+r_{1}\eta)+k_{2}q_{0}$. Hence,
$\inf_{r_{1}\in[0,1)}[G(s+r_{1}\eta)-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(t+r_{2}\eta)-G(r_{2}\eta)]$
$=\inf_{r_{1}\in[0,1)}[G(s-t+r_{1}\eta)+k_{2}q_{0}-G(r_{1}\eta)]-k_{2}q_{0}$
$=\inf_{r_{1}\in[0,1)}[G(s-t+r_{1}\eta)-G(r_{1}\eta)]\geq\inf_{r_{1}\in[0,1]}[G(s-t+r_{1}\eta)-G(r_{1}\eta)].$
Since $r_{1}\mapsto G(s-t+r_{1}\eta)-G(r_{1}\eta)$ is a continuous function and $[0,1]$ is a compact set, the above infimum is achieved at point in $[0,1]$. Since $G$ is strictly increasing and $s-t>0$, we have
$\Delta_{2}:=\inf_{r_{1}\in[0,1]}[G(s-t+r_{1}\eta)-G(r_{1}\eta)]>0.$
We now choose $\varepsilon\in(0,(s-t)/4]$ such that
$\sup_{r_{2}\in[0,1)}[G(\varepsilon+r_{2}\eta)-G(r_{2}\eta)]\leq\Delta_{2}/4$
and
$\sup_{r_{1}\in[0,1]}[G(s-t+r_{1}\eta)-G(s-\varepsilon-t+r_{1}\eta)]\leq\Delta_{2}/4$
To see that this is possible, again notice that $G(\cdot)$ is uniformly continuous on the compact set $[0,\eta+1]\bigcup[(s-t),\eta+1+(s-t)]$. Thus, there exists a constant $\kappa_{2}>0$ such that $\sup_{r_{2}\in[0,1)}[G(\varepsilon+r_{2}\eta)-G(r_{2}\eta)]\leq\kappa_{2}\varepsilon$ and $\sup_{r_{1}\in[0,1]}[G(s-t+r_{1}\eta)-G(s-\varepsilon-t+r_{1}\eta)]\leq\kappa_{2}\varepsilon$. Hence, we can simply choose $\varepsilon=\min\{(s-t)/4,\Delta_{2}/(4\kappa_{2})\}$. Therefore,
$\inf_{r_{1}\in[0,1)}[G(s-\varepsilon+r_{1}\eta)-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(t+\varepsilon+r_{2}\eta)-G(r_{2}\eta)]$
$=\inf_{r_{1}\in[0,1)}[G(s-\varepsilon-t+r_{1}\eta)+k_{2}q_{0}-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(\varepsilon+r_{2}\eta)+k_{2}q_{0}-G(r_{2}\eta)]$
$=\inf_{r_{1}\in[0,1)}[G(s-\varepsilon-t+r_{1}\eta)-G(r_{1}\eta)]-\sup_{r_{2}\in[0,1)}[G(\varepsilon+r_{2}\eta)-G(r_{2}\eta)]$
$\geq\inf_{r_{1}\in[0,1)}[G(s-\varepsilon-t+r_{1}\eta)-G(s-t+r_{1}\eta)]+\inf_{r_{1}\in[0,1)}[G(s-t+r_{1}\eta)-G(r_{1}\eta)]$
$\quad-\sup_{r_{2}\in[0,1)}[G(\varepsilon+r_{2}\eta)-G(r_{2}\eta)]$
$=-\sup_{r_{1}\in[0,1)}[G(s-t+r_{1}\eta)-G(s-\varepsilon-t+r_{1}\eta)]+\inf_{r_{1}\in[0,1)}[G(s-t+r_{1}\eta)-G(r_{1}\eta)]$
$\quad-\sup_{r_{2}\in[0,1)}[G(\varepsilon+r_{2}\eta)-G(r_{2}\eta)]$
[…]
where (i) follows by the construction of $\varepsilon$ and (ii) follows by the definition of $\Delta_{2}$.
Thus, we have verified (22) and thus, $\inf_{\alpha\in\mathbb{R}}[G(s-\varepsilon+\alpha)-G(\alpha)]>\sup_{\xi\in\mathbb{R}}[G(t+\varepsilon+\xi)-G(\xi)]$ for some $\varepsilon\in(0,(s-t)/4]$. Clearly $\sup_{\xi\in\mathbb{R}}[G(t+\varepsilon+\xi)-G(\xi)]\geq G(t+\varepsilon)-G(0)>0$ since $t,\varepsilon>0$ and $G$ is strictly increasing. ∎
Lemma 14.
Let $\eta>0$. If $\mathcal{Z}$ is compact and has non-empty interior, then there exists an open set $D\subset\mathcal{B}_{+}$ such that for any $d=(d_{1}^{\prime},d_{2})^{\prime}\in D$, there exists $z\in\mathcal{Z}^{\circ}$ such that $(d_{1}^{\prime}z+d_{2})/\eta$ is an integer.
Proof.
Let $R=\sup_{z\in\mathcal{Z}}\|z\|_{2}$. Since $\mathcal{Z}$ is compact, $R<\infty$. Fix any $q=(q_{1}^{\prime},q_{2})^{\prime}\in\mathbb{R}^{K}$ with $q_{2}=\|q_{1}\|_{2}R+1$ and $q_{1}\neq 0$. Fix an $z_{0}\in\mathcal{Z}^{\circ}$. Then there exists $\varepsilon>0$ such that $\mathbb{B}(z_{0},\varepsilon)\subset\mathcal{Z}^{\circ}$, where $\mathbb{B}(z_{0},\varepsilon):=\{z:\ \|z-z_{0}\|_{2}\leq\varepsilon\}$. Define $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}$ with $\beta_{1}=\alpha q_{1}$, $\beta_{2}=\alpha q_{2}$ and $\alpha=\eta/(z_{0}^{\prime}q_{1}+q_{2})$. Clearly, $z^{\prime}q_{1}+q_{2}\geq-R\|q_{1}\|+q_{2}=1>0$ for any $z\in\mathcal{Z}$. Thus, $\alpha>0$ and $z^{\prime}\beta_{1}+\beta_{2}\geq\alpha$ for any $z\in\mathcal{Z}$. Hence, $\beta\in\mathcal{B}_{+}$ and $z_{0}^{\prime}\beta_{1}+\beta_{2}=\eta$.
Fix any $\delta>0$ such that $\delta\leq\alpha\|q_{1}\|_{2}/(1+\sqrt{\|z_{0}\|_{2}^{2}+1})$ and $\delta\leq\alpha/[2\sqrt{R^{2}+1}]$.
Now consider any $d=(d_{1}^{\prime},d_{2})^{\prime}$ such that $\|d-\beta\|_{2}\leq\delta$. We notice that for any $z\in\mathcal{Z}$, $|z^{\prime}(d_{1}-\beta_{1})+(d_{2}-\beta_{2})|\leq\sqrt{\|z\|_{2}^{2}+1}\|d-\beta\|_{2}\leq\sqrt{R^{2}+1}\delta\leq\alpha/2$. Thus, for any $z\in\mathcal{Z}$, $z^{\prime}d_{1}+d_{2}\geq z^{\prime}\beta_{1}+\beta_{2}-\alpha/2\geq\alpha-\alpha/2>0$. Thus, $d\in\mathcal{B}_{+}$.
We observe that
$\max_{z\in\mathbb{B}(z_{0},\varepsilon)}d_{1}^{\prime}z+d_{2}=d_{1}^{\prime}z_{0}+d_{2}+\varepsilon\max_{\|v\|_{2}\leq 1}d_{1}^{\prime}v=d_{1}^{\prime}z_{0}+d_{2}+\varepsilon\|d_{1}\|_{2}$
and
$\min_{z\in\mathbb{B}(z_{0},\varepsilon)}d_{1}^{\prime}z+d_{2}=d_{1}^{\prime}z_{0}+d_{2}+\varepsilon\min_{\|v\|_{2}\leq 1}d_{1}^{\prime}v=d_{1}^{\prime}z_{0}+d_{2}-\varepsilon\|d_{1}\|_{2}.$
Notice that the maximum and minimum over $\mathbb{B}(z_{0},\varepsilon)$ are achieved (since $\mathbb{B}(z_{0},\varepsilon)$ is compact). Hence, by the intermediate value theorem and the convexity of $\mathbb{B}(z_{0},\varepsilon)$, it suffices to show that $d_{1}^{\prime}z_{0}+d_{2}-\varepsilon\|d_{1}\|_{2}\leq\eta\leq d_{1}^{\prime}z_{0}+d_{2}+\varepsilon\|d_{1}\|_{2}$. In other words, it suffices to show that
$|d_{1}^{\prime}z_{0}+d_{2}-\eta|\leq\varepsilon\|d_{1}\|_{2}.$
By $\eta=z_{0}^{\prime}\beta_{1}+\beta_{2}$, we have $|d_{1}^{\prime}z_{0}+d_{2}-\eta|=|(d_{1}-\beta_{1})^{\prime}z_{0}+d_{2}-\beta_{2}|\leq\|(z_{0}^{\prime},1)\|_{2}\|d-\beta\|_{2}\leq\sqrt{\|z_{0}\|_{2}^{2}+1}\delta$ and $\|d_{1}\|_{2}\geq\|\beta_{1}\|_{2}-\|\beta_{1}-d_{1}\|_{2}\geq\|\beta_{1}\|_{2}-\delta=\alpha\|q_{1}\|_{2}-\delta$. By construction, $\sqrt{\|z_{0}\|_{2}^{2}+1}\delta\leq\alpha\|q_{1}\|_{2}-\delta$. The above display holds. Hence, the result holds with $D=\{d:\|d-\beta\|_{2}<\delta\}$. ∎
Proof of Theorem 4.
The first claim follows by Lemma 11. The third claim follows by Lemma 14 and the second claim. It remains to prove the second claim.
We partition $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}$. By assumption, there exists $z_{0}\in\mathcal{Z}^{\circ}$ such that $k=(z_{0}^{\prime}\beta_{1}+\beta_{2})/\eta$ is an integer. Clearly $k>0$ since $\beta\in\mathcal{B}_{+}$ (i.e., $z^{\prime}\beta_{1}+\beta_{2}>0$ for any $z\in\mathcal{Z}$). Consider any $b\in\mathcal{B}_{+}$ such that $b=(b_{1}^{\prime},b_{2})^{\prime}\neq\beta$. We proceed in two steps.
Step 1: show that there exists $z_{*}\in\mathcal{Z}^{\circ}$ such that $(z_{*}^{\prime},1)\beta\neq(z_{*}^{\prime},1)b$ and one of $(z_{*}^{\prime},1)\beta/\eta$ and $(z_{*}^{\prime},1)b/\eta$ is an integer.
Suppose that $z_{0}^{\prime}b_{1}+b_{2}\neq k\eta$. Then we take $z_{*}=z_{0}$ and have $(z_{*}^{\prime},1)\beta=k\eta\neq(z_{*}^{\prime},1)b$. Clearly, one of $(z_{*}^{\prime},1)\beta/\eta$ and $(z_{*}^{\prime},1)b/\eta$ is an integer.
Suppose that $z_{0}^{\prime}b_{1}+b_{2}=k\eta$. We notice that $\beta_{1}\neq b_{1}$. To see this, suppose $\beta_{1}=b_{1}$. Then $z_{0}^{\prime}b_{1}=z_{0}^{\prime}\beta_{1}$. Notice that $z_{0}^{\prime}b_{1}+b_{2}=k\eta=z_{0}^{\prime}\beta_{1}+\beta_{2}$. We have $b_{2}=\beta_{2}$. By $\beta_{1}=b_{1}$, this would imply $\beta=b$, contradicting $\beta\neq b$. Hence, $\beta_{1}\neq b_{1}$ and thus at least one of $\beta_{1}$ and $b_{1}$ is not zero. Then we only need to discuss three cases.
Case 1: suppose $\beta_{1}\neq 0$ and $b_{1}\neq 0$. Since $z_{0}\in\mathcal{Z}^{\circ}$ and $\beta_{1}\neq 0$, $z_{1}=z_{0}+(I-\beta_{1}\beta_{1}/\|\beta_{1}\|_{2}^{2})\Delta\in\mathcal{Z}^{\circ}$ when $\|\Delta\|_{2}$ is small enough. We choose $\Delta$ such that $b_{1}^{\prime}(I-\beta_{1}\beta_{1}/\|\beta_{1}\|_{2}^{2})\Delta>0$; this is possible because $b_{1}\neq\beta_{1}$, $\beta_{1}\neq 0$ and $b_{1}\neq 0$. Then $z_{1}^{\prime}\beta_{1}+\beta_{2}=z_{0}^{\prime}\beta_{1}+\beta_{2}=k\eta$ and $z_{1}^{\prime}b_{1}+b_{2}=z_{0}^{\prime}b_{1}+b_{2}+(z_{1}-z_{0})^{\prime}b_{1}=k\eta+b_{1}^{\prime}(I-\beta_{1}\beta_{1}/\|\beta_{1}\|_{2}^{2})\Delta>k\eta$. We now define $t=z_{1}^{\prime}\beta_{1}+\beta_{2}$ and $s=z_{1}^{\prime}b_{1}+b_{2}$. Thus $s=k\eta>t>0$. In other words, we can take $z_{*}=z_{1}$.
Case 2: suppose $\beta_{1}\neq 0$ and $b_{1}=0$. Since $z_{0}^{\prime}b_{1}+b_{2}=k\eta$ and $b_{1}=0$, we have $b_{2}=k\eta$. Since $\mathcal{Z}^{\circ}$ contains an open neighborhood of $z_{0}$ and $\beta_{1}\neq 0$, we can find $z_{1}$ close enough to $z_{0}$ such that $(z_{1}-z_{0})^{\prime}\beta_{1}>0$. This means that $z_{1}^{\prime}\beta_{1}+\beta_{2}=(z_{1}-z_{0})^{\prime}\beta_{1}+(z_{0}^{\prime}\beta_{1}+\beta_{2})>k\eta$ and $z_{1}^{\prime}b_{1}+b_{2}=b_{2}=k\eta$. We now define $s=z_{1}^{\prime}\beta_{1}+\beta_{2}$ and $t=z_{1}^{\prime}b_{1}+b_{2}$. Thus, $s>t=k\eta>0$. In other words, we can take $z_{*}=z_{1}$.
Case 3: suppose $\beta_{1}=0$ and $b_{1}\neq 0$. The argument is analogous to the above case. We repeat it here for completeness. Since $z_{0}^{\prime}\beta_{1}+\beta_{2}=k\eta$ and $\beta_{1}=0$, we have $\beta_{2}=k\eta$. Since $\mathcal{Z}^{\circ}$ contains an open neighborhood of $z_{0}$ and $b_{1}\neq 0$, we can find $z_{1}$ close enough to $z_{0}$ such that $(z_{1}-z_{0})^{\prime}b_{1}>0$. This means that $z_{1}^{\prime}b_{1}+b_{2}=(z_{1}-z_{0})^{\prime}b_{1}+(z_{0}^{\prime}b_{1}+b_{2})>k\eta$ and $z_{1}^{\prime}\beta_{1}+\beta_{2}=\beta_{2}=k\eta$. We now define $s=z_{1}^{\prime}b_{1}+b_{2}$ and $t=z_{1}^{\prime}\beta_{1}+\beta_{2}$. Thus, $s>t=k\eta>0$. In other words, we can take $z_{*}=z_{1}$.
Step 2: show the final result.
We now prove the result in two cases.
Case 1: suppose $(z_{*}^{\prime},1)\beta>(z_{*}^{\prime},1)b$. Define $s=(z_{*}^{\prime},1)^{\prime}\beta$ and $t=(z_{*}^{\prime},1)^{\prime}b$. Then $s>t>0$ and either $s/\eta$ or $t/\eta$ is an integer. By Lemma 13, there exists $\varepsilon\in(0,(s-t)/4]$ such that
$\inf_{\alpha\in\mathbb{R}}[G(s-\varepsilon+\alpha)-G(\alpha)]>\sup_{\xi\in\mathbb{R}}[G(t+\varepsilon+\xi)-G(\xi)]>0.$
Let $\mathbb{B}=\{v\in\mathbb{R}^{K-1}:\|v\|_{2}<1\}$ be the open unit ball. Notice that $z_{*}\in\mathcal{Z}^{\circ}$. Clearly, there exists $\tau>0$ such that $z_{*}+\tau\mathbb{B}\subset\mathcal{Z}^{\circ}$, $\tau|v^{\prime}\beta_{1}|\leq\varepsilon$ and $\tau|v^{\prime}b_{1}|\leq\varepsilon$ for any $v\in\mathbb{B}$. Define $\mathcal{Z}_{\tau}=z_{*}+\tau\mathbb{B}$. By construction, $\mathcal{Z}_{\tau}$ is an open set contained in $\mathcal{Z}^{\circ}$. Notice that for any $z\in\mathcal{Z}_{\tau}$, $(z^{\prime},1)^{\prime}\beta=(z_{*}^{\prime},1)^{\prime}\beta+(z-z_{*})^{\prime}\beta_{1}\geq(z_{*}^{\prime},1)^{\prime}\beta-\varepsilon=s-\varepsilon$ and $(z^{\prime},1)^{\prime}b=(z_{*}^{\prime},1)^{\prime}b+(z-z_{*})^{\prime}\beta_{1}\leq(z_{*}^{\prime},1)^{\prime}b+\varepsilon=t+\varepsilon$. Thus, the above display implies that for any $z\in Z_{\tau}$, $(z^{\prime},1)^{\prime}\beta\geq s-\varepsilon>t+\varepsilon\geq(z^{\prime},1)^{\prime}b$ and thus
$\inf_{\alpha\in\mathbb{R}}[G((z^{\prime},1)^{\prime}\beta+\alpha)-G(\alpha)]\geq\inf_{\alpha\in\mathbb{R}}[G(s-\varepsilon+\alpha)-G(\alpha)]\\ >\sup_{\xi\in\mathbb{R}}[G(t+\varepsilon+\xi)-G(\xi)]\geq\sup_{\xi\in\mathbb{R}}[G((z^{\prime},1)^{\prime}b+\xi)-G(\xi)]>0.$
By Lemma 12, it follows that for any $z\in\mathcal{Z}_{\tau}$, $\mathcal{A}((z^{\prime},1)^{\prime}\beta)\bigcap\mathcal{A}((z^{\prime},1)^{\prime}b)=\emptyset$.
Case 2: suppose $(z_{*}^{\prime},1)\beta<(z_{*}^{\prime},1)b$. The argument is analogous to Case 1 except that we swap the roles of $\beta$ and $b$. We omit it for simplicity.
Therefore, we have proved that for any $b\in\mathcal{B}_{+}$ with $b\neq\beta$, there exists an open set $\mathcal{Z}_{b}\subset\mathcal{Z}$ such that for any $z\in\mathcal{Z}_{b}$, $\mathcal{A}(w^{\prime}\beta)\bigcap\mathcal{A}(w^{\prime}b)=\emptyset$ with $w=(z^{\prime},1)^{\prime}$. Clearly $\mathcal{Z}_{b}\bigcap\mathcal{Z}\neq\emptyset$. By Lemma 3, $P(Z\in\mathcal{Z}_{b})>0$. ∎
A.6 Proof of Theorem 5
In the proof of Theorem 3, we proved that any $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}\in\mathcal{B}_{+}$ is observationally equivalent to $b=(\beta_{1}^{\prime},\beta_{2}+\delta)^{\prime}$ for some $\delta>0$. Now we pick any $\beta\in\mathcal{B}_{+,0}$. Then $b=(\beta_{1}^{\prime},\beta_{2}+\delta)^{\prime}=(\beta_{1}^{\prime},\delta)\in\mathcal{B}_{+,+}$. The desired result follows.
A.7 Proof of results in Section 4
Proof of Lemma 2.
Notice that $\rho(q)=E\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})=E\mathbf{1}\{W^{\prime}q\geq 0\}\cdot\phi(X)$. Then by construction $\rho(q)\leq E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)$ for any $q$, which means
$\sup_{q\in\mathbb{R}^{K}}\rho(q)\leq E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X).$
On the other hand,
[…]
where (i) follows by Lemma 1. The above two displays imply $\sup_{q\in\mathbb{R}^{K}}\rho(q)=E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)=E\max\{\phi(X),0\}$. The proof of the other claim follows analogously with $\inf_{q}\rho(q)=\rho(-\beta)=E\min\{\phi(X),0\}$. ∎
Proof of Theorem 6.
Define $S_{n}(q)=\sqrt{n}(\hat{\rho}_{n}(q)-\rho(q))$. Notice that the class of mappings $W\mapsto\mathbf{1}\{W^{\prime}q\geq 0\}$ indexed by $q$ has VC-dimension at most $K+2$ (Lemmas 2.6.15 and 2.6.18 of van der Vaart and Wellner, (1996)). Clearly, the class $\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})$ indexed by $q$ is bounded. By Theorem 2.5.2 of van der Vaart and Wellner, (1996), $S_{n}(\cdot)$ converges weakly to a mean-zero Gaussian process $S_{*}(\cdot)$. By Theorem 3.6.1 of van der Vaart and Wellner, (1996), we can approximate the distribution of $S_{*}$ via the nonparametric bootstrap.
Of course, we need to check that $S_{*}$ is not degenerate under $H_{0}$. By Lemma 1, under $H_{0}$, $P(W^{\prime}\beta\leq 0)=1$,
$\rho(-\beta)=E\mathbf{1}\{W^{\prime}\beta\leq 0\}(Y_{1}-Y_{0})=E(Y_{1}-Y_{0})$
and thus
$E(S_{n}(-\beta))^{2}$ $=E\mathbf{1}\{W^{\prime}\beta\leq 0\}(Y_{1}-Y_{0})^{2}-(\rho(\beta))^{2}$
$=E(Y_{1}-Y_{0})^{2}-\left(E(Y_{1}-Y_{0})\right)^{2}={\rm Var}(Y_{1}-Y_{0}).$
Therefore, $\mathbf{1}\{W^{\prime}\beta<0\}(Y_{1}-Y_{0})$ is not degenerate since $Y_{1}-Y_{0}$ is non-zero variance (due to the assumption of $P(Y_{1}=Y_{0})<1$).
Let $c_{1-\alpha}^{*}$ be such that $\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}S_{*}(q)>c_{1-\alpha}^{*}\right)=\alpha$. If $\tau_{1}\leq-\tau_{2}$, then
$\sqrt{n}\left(\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-\min\{\tau_{1},-\tau_{2}\}\right)$
$=\sqrt{n}\left(\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-\tau_{1}\right)$
$=\sqrt{n}\min\{\hat{\tau}_{1}-\tau_{1},\tau_{1}-\hat{\tau}_{2}\}$
$\leq\sqrt{n}(\hat{\tau}_{1}-\tau_{1})=\sqrt{n}\left(\sup_{q\in\mathbb{R}^{K}}\hat{\rho}(q)-\sup_{q\in\mathbb{R}^{K}}\rho(q)\right)\leq\sqrt{n}\sup_{q\in\mathbb{R}^{K}}(\hat{\rho}(q)-\rho(q))=\sup_{q\in\mathbb{R}^{K}}S_{n}(q).$
If $\tau_{1}>-\tau_{2}$, then
$\sqrt{n}\left(\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-\min\{\tau_{1},-\tau_{2}\}\right)$
$=\sqrt{n}\left(\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}+\tau_{2}\right)$
$=\sqrt{n}\min\{\hat{\tau}_{1}+\tau_{2},\tau_{2}-\hat{\tau}_{2}\}$
$\leq\sqrt{n}(\tau_{2}-\hat{\tau}_{2})=\sqrt{n}\left(\inf_{q\in\mathbb{R}^{K}}\rho(q)-\inf_{q\in\mathbb{R}^{K}}\hat{\rho}(q)\right)\leq\sqrt{n}\sup_{q\in\mathbb{R}^{K}}(\rho(q)-\hat{\rho}(q))=\sup_{q\in\mathbb{R}^{K}}\left(-S_{n}(q)\right).$
Therefore,
$\sqrt{n}\left(\hat{\tau}_{*}-\tau_{*}\right)\leq\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)\right)\cdot\mathbf{1}\{\tau_{1}\leq-\tau_{2}\}+\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))\right)\cdot\mathbf{1}\{\tau_{1}>-\tau_{2}\}.$
Therefore,
$P\left(\sqrt{n}\left(\hat{\tau}_{*}-\tau_{*}\right)>c_{1-\alpha}^{*}\right)\\ \leq P\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)>c_{1-\alpha}^{*}\right)\cdot\mathbf{1}\{\tau_{1}\leq-\tau_{2}\}+P\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))>c_{1-\alpha}^{*}\right)\cdot\mathbf{1}\{\tau_{1}>-\tau_{2}\}.$
Since $S_{n}(\cdot)$ converges weakly to $S_{*}(\cdot)$, $P\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)>c_{1-\alpha}^{*}\right)$ and $P\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))>c_{1-\alpha}^{*}\right)$ converge to $\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}(S_{*}(q))>c_{1-\alpha}^{*}\right)$ and $\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}(-S_{*}(q))>c_{1-\alpha}^{*}\right)$, respectively. Both limits are equal to $\alpha$; this is because $-S_{*}(\cdot)$ and $S_{*}(\cdot)$ have the same distribution. The result follows since we already proved that $S_{*}(\cdot)$ can be bootstrapped.∎